### v1

# Non-commutative Iwasawa theory of abelian varieties over global function fields

Abstract.
Let $A$ be an abelian variety defined over a global function field $F$. We investigate the structure of the $p$-primary Selmer group $\mathrm{Sel}(A/F_{\infty})$ for any prime number $p$ distinct from the characteristic of $F$, over $p$-adic Lie extensions $F_{\infty}$ of $F$ which contain the cyclotomic $\mathbb{Z}_{p}$-extension $F^{\mathrm{cyc}}$. In particular, we prove that the Pontryagin dual of the Selmer group $\mathrm{Sel}(A/F^{\mathrm{cyc}})$ is a torsion $\mathbb{Z}_{p}[[\mathrm{Gal}(F^{\mathrm{cyc}}/F)]]$-module with trivial $\mu$-invariant, and we establish the $\mathfrak{M}_{H}(G)$-conjecture of Coates–Fukaya–Kato–Sujatha–Venjakob for $A/F_{\infty}$.
In view of the validity of the $\mathfrak{M}_{H}(G)$-conjecture, it therefore makes sense to speak of the characteristic element (in the sense of Coates et al.) attached to the Pontryagin dual of $\mathrm{Sel}(A/F_{\infty})$. We relate the order of vanishing of the characteristic elements, evaluated at Artin representations, to the corank of the Selmer group of the corresponding twist of $A$ over the base field $F$. Combining this with the deep results of Tate, Milne and Kato–Trihan, we show that the order of vanishing of the characteristic elements is equal to the order of vanishing of the $L$-function of $A/F$ at $s=1$ under appropriate assumptions.
Finally, we relate the generalised Euler characteristic of $\mathrm{Sel}(A/F_{\infty})$ to the Euler characteristic of $\mathrm{Sel}(A/F^{\mathrm{cyc}})$. This is a natural analogue of Zerbes’ result in the number field context and generalises previous results of Sechi and Valentino in the function field context.
Key words and phrases:
Elliptic curves, Iwasawa theory, Euler characteristics
2010 Mathematics Subject Classification:
11R23 (primary), 11R34 (secondary)
The second author is supported by the ANR project Coloss and the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 101026826.
1. Introduction
Non-commutative Iwasawa theory has emerged as a powerful framework for understanding deep arithmetic properties over number fields contained in a $p$-adic Lie extension and their precise relationship to special values of complex $L$-functions. The aim of the present paper is to study non-commutative Iwasawa theory over global function fields. To motivate our discussion, we begin by reviewing the number field context as developed by Coates, Fukaya, Kato, Sujatha, and Venjakob in [8] (see also [3, 4, 38]). In the number field case, $p$ will denote an odd prime. Given a compact $p$-adic Lie group $\mathcal{G}$, the Iwasawa algebra $\Lambda(\mathcal{G})$ of $\mathcal{G}$ is defined to be the completed group ring
[…]
where $\mathcal{U}$ runs over all open normal subgroups of $\mathcal{G}$. Let $E$ be an elliptic curve defined over a number field $F$ with good ordinary reduction at all primes of $F$ above $p$. Denote by $F_{\infty}$ an admissible $p$-adic Lie extension of $F$, that is, a Galois extension of $F$ such that $F^{\mathrm{cyc}}\subset F_{\infty}$ and whose Galois group $G={\mathrm{Gal}}(F_{\infty}/F)$ is a $p$-adic Lie group of positive dimension. Here $F^{\mathrm{cyc}}$ denotes the cyclotomic $\mathbb{Z}_{p}$-extension of $F$. Write $\Gamma={\mathrm{Gal}}(F^{\mathrm{cyc}}/F)$, $H={\mathrm{Gal}}(F_{\infty}/F^{\mathrm{cyc}})$, and set
$X(E/F_{\infty})={\mathrm{Hom}}_{\rm cts}({\mathrm{Sel}}(E/F_{\infty}),{\mathbb{Q}}_{p}/{\mathbb{Z}}_{p})$
for the Pontryagin dual of the $p$-primary Selmer group ${\mathrm{Sel}}(E/F_{\infty})$, endowed with its natural left $\Lambda(G)$-module structure. In the case where $F_{\infty}=F^{\mathrm{cyc}}$, it is well-known that $X(E/F^{\mathrm{cyc}})$ is a finitely generated $\Lambda(\Gamma)$-module. It is generally expected that $X(E/F^{\mathrm{cyc}})$ enjoys nice module-theoretical properties, as illustrated by the following conjecture of Mazur [24]:
Conjecture 1.1 (Mazur).
The Iwasawa module $X(E/F^{\mathrm{cyc}})$ is a finitely generated torsion $\Lambda(\Gamma)$-module.
This has been verified by Mazur himself in the case where ${\mathrm{Sel}}(E/F)$ is finite. The best evidence in support of this conjecture is a deep result of Kato [17], where the torsionness assertion is established when $E$ is defined over ${\mathbb{Q}}$ and $F$ is an abelian extension of ${\mathbb{Q}}$. Taking the structure theory of $\Lambda(\Gamma)$-module into account, Mazur’s conjecture is equivalent to saying that $X(E/F^{\mathrm{cyc}})/X(E/F^{\mathrm{cyc}})(p)$ is finitely generated over ${\mathbb{Z}}_{p}$, where $X(E/F^{\mathrm{cyc}})(p)$ denotes the $p$-primary submodule of $X(E/F^{\mathrm{cyc}})$. Motivated by this observation, Coates et al. introduce the category ${\mathfrak{M}}_{H}(G)$ which consists of all finitely generated $\Lambda(G)$-modules $M$ whose quotient $M/M(p)$ by the $p$-primary submodule $M(p)$ is a finitely generated $\Lambda(H)$-module. They then conjectured that ${\mathfrak{M}}_{H}(G)$ contains all the torsion $\Lambda(G)$-modules which are of interest in arithmetic applications. More precisely, they formulated (see Section 5 and Conjecture 5.1 of [8]):
Conjecture 1.2 (${\mathfrak{M}}_{H}(G)$-Conjecture).
The Iwasawa module $X(E/F_{\infty})$ belongs to the category ${\mathfrak{M}}_{H}(G)$.
This conjecture is crucial as it allows one to define a characteristic element for $X(E/F_{\infty})$. This is essential for formulating the main conjectures of non-commutative Iwasawa theory (see [8, 4]). Coates et al. further demonstrated in their paper how these characteristic elements relate to the Euler characteristics of $X(E/F_{\infty})$ via the theory of Akashi series introduced in [9]. This latter aspect has been further explored by Zerbes in a series of papers [39, 40, 41], where she related the characteristic elements to the generalised Euler characteristics of $X(E/F_{\infty})$. Here, the notion of the generalised Euler characteristics is an extension of the usual Euler characteristics, whose idea emerged in [9], and was later developed and systematically studied by Zerbes in [40].
We now shift our focus to the function field setting, the central theme of our paper. Henceforth, $p$ denotes a fixed prime, where we allow $p=2$. Throughout this paper, $F$ will always denote a global function field. In other words, $F$ is a finite extension of the field $k(T)$ of rational functions in one variable over a finite constant field $k$. We always assume that the characteristic $\mathrm{char}(F)$ of $F$ is distinct from our fixed prime $p$. We fix a separable closure $\bar{F}$ of $F$. The cyclotomic ${\mathbb{Z}}_{p}$-extension $F^{\rm cyc}$ of $F$ is obtained by adjoining to $F$ the unique ${\mathbb{Z}}_{p}$-constant field extension of $k$. Thus, writing $\mu_{p^{\infty}}$ for the group of $p$-power roots of unity in an algebraic closure $\bar{k}$ of $k$, $F^{\rm cyc}$ is the unique ${\mathbb{Z}}_{p}$-extension of $F$ contained in $F(\mu_{p^{\infty}})$. Following the analogy with the number field setting, we define an admissible $p$-adic Lie extension of $F$ as an extension $F_{\infty}$ satisfying the following conditions:
  1. (i)
    $F_{\infty}$ contains the cyclotomic ${\mathbb{Z}}_{p}$-extension $F^{\mathrm{cyc}}$ of $F$.
  2. (ii)
    $G={\mathrm{Gal}}(F_{\infty}/F)$ is a compact $p$-adic group without $p$-torsion.
  3. (iii)
    $F_{\infty}/F$ is unramified outside of a finite set of primes.
It is worth noting that a global function field $F$ with $\mathrm{char}(F)\neq p$ has only one ${\mathbb{Z}}_{p}$-extension, the cyclotomic one (see [28, Proposition 10.3.20]). Therefore, any admissible extension of dimension at least $2$ is necessarily non-commutative. This underscores the relevance of non-commutative Iwasawa theory in the context of function fields. We also note that when $\mathrm{char}(F)=p$, there are infinitely many geometric (i.e., not containing any constant field extension) ${\mathbb{Z}}_{p}$-extensions of $F$ (see [14, Theorem 3]). This presents another intriguing difference between number fields and function fields.
We now return to the core of the paper. Our first main result is as follows:
Theorem 1.3.
Let $A$ be an abelian variety defined over a global function field $F$ with characteristic prime to $p$. Let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Then, the $\mathfrak{M}_{H}(G)$-conjecture holds for the Pontryagin dual $X(A/F_{\infty})$ of the Selmer group ${\mathrm{Sel}}(A/F_{\infty})$.
Furthermore, the Pontryagin dual $X(A/F^{\mathrm{cyc}})$ of the Selmer group ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ has $\mu$-invariant equal to zero.
In particular, our result completely resolves Mazur’s conjecture and the $\mathfrak{M}_{H}(G)$-conjecture in the function field setting. This presents a stark contrast with the number field case. It is long established that the Pontryagin dual of the Selmer group over the cyclotomic ${\mathbb{Z}}_{p}$-extension of a number field can have a positive $\mu$-invariant (see [24, §10, Example 2], [13]). In contrast, our result shows that this never occurs in function fields. We note that Sechi [31] has given a proof of this theorem for elliptic curves $E$ without complex multiplication defined over $F$, where $F_{\infty}$ is obtained by adjoining the $p$-power torsion points of $E$ to $F$, under the assumption that ${\mathrm{Sel}}(E/F)$ is finite. In the case $\mathrm{char}(F)=p$, the ${\mathfrak{M}}_{H}(G)$-conjecture has been proven by Ochiai and Trihan under a $\mu=0$ assumption (see [29, Theorem 1.9]). Our proof of this theorem differs entirely from previous work on the ${\mathfrak{M}}_{H}(G)$-conjecture. In particular, Theorem 1.3 provides the first example where the ${\mathfrak{M}}_{H}(G)$-conjecture holds unconditionally. A key element of our proof is the observation that the cohomological dimension of ${\mathrm{Gal}}(\bar{F}/F^{\mathrm{cyc}})$ is equal to one. From there, our approach is essentially cohomological in nature. The authors find it curious that this line of approach seems to have been overlooked in literature.
In light of this result on the $\mathfrak{M}_{H}(G)$-conjecture, we can now attach a characteristic element $\xi_{A}$ to $X(A/F_{\infty})$ in the sense of Coates et al. This allows us to evaluate $\xi_{A}$ at a given Artin representation $\rho:G\to{\mathrm{GL}}_{n}(\mathcal{O})$, where $\mathcal{O}$ is the ring of integers of a finite extension of ${\mathbb{Q}}_{p}$ in ¯$\overline{{\mathbb{Q}}}_{p}$, via the homomorphism $\Phi_{\rho}$ (for details, see Section 7, in particular (7.2)). This yields an element in the field of fractions of the formal power series ring $\mathcal{O}[[T]]$ in an indeterminate $T$ with coefficients in $\mathcal{O}$. For subsequent discussion, we shall introduce the following hypothesis to facilitate explicit calculations:
Hypothesis ($\mathbf{G}$). $H^{i}(G,A_{p^{\infty}}(F_{\infty}))$ is finite for all $i\geq 1$.
We are now in a position to state our next result, presented here in a slightly simplified form. For a more precise and general statement of the theorem below, we refer readers to Theorem 7.7, Corollary 7.8, and Theorem 7.12.
Theorem 1.4.
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Denote by $\xi_{A}$ the characteristic element of $X(A/F_{\infty})$, and let $\mathrm{reg}_{F}$ denote the regular representation of $F$. Assuming Hypothesis $(\mathbf{G})$ holds, we have
${\mathrm{ord}}_{T=0}(\Phi_{\mathrm{reg}_{F}}(\xi_{A}))\geq{\mathrm{corank}}_{{\mathbb{Z}}_{p}}({\mathrm{Sel}}(A/F)).$
In addition, equality holds when Greenberg’s semi-simplicity conjecture 6.2 holds for $A/F$. Moreover, if we assume the finiteness of $\hbox{\cyr X}(A/F)$, then
${\mathrm{ord}}_{T=0}(\Phi_{\mathrm{reg}_{F}}(\xi_{A}))={\mathrm{ord}}_{s=1}L(A/F,s).$
In fact, the finiteness assumption on the $\ell$-primary part of the Tate-Shafarevich group $\hbox{\cyr X}(A/F)(\ell)$ for a single prime number $\ell$ suffices for the last part of the above theorem by the deep theorem of Kato–Trihan [18].
In order to state our final result, we need to introduce the following stronger finiteness hypothesis. Recall that we write $H={\mathrm{Gal}}(F_{\infty}/F^{\rm cyc})$.
Hypothesis ($\mathbf{H}$). $H^{i}(H,A_{p^{\infty}}(F_{\infty}))$ is finite for all $i\geq 0$.
Let $S$ be a nonempty finite set of primes of $F$ containing all the primes where the abelian variety $A$ has bad reduction, and all the primes of $F$ which ramify in $F_{\infty}$, and let $S^{\prime}$ denotes the subset of primes whose inertia group in $G$ is infinite. We write $A^{*}$ for the dual abelian variety. The following is our final main result.
Theorem 1.5.
Let $A$ be an abelian variety defined over a global field $F$ with $\mathrm{char}F\neq p$ and let $F_{\infty}$ be an admissible $p$-adic Lie extension of $F$. Assume that Hypothesis $(\mathbf{H})$ holds. Then ${\mathrm{Sel}}(A/F_{\infty})$ has finite generalised $G$-Euler characteristic $\chi(G,{\mathrm{Sel}}(A/F_{\infty}))$ if and only if ${\mathrm{Sel}}(A/F^{\rm cyc})$ has finite generalised $\Gamma$-Euler characteristic $\chi(\Gamma,{\mathrm{Sel}}(A/F^{\mathrm{cyc}}))$. Moreover, if this is the case, then
$\chi(G,{\mathrm{Sel}}(A/F_{\infty}))=\chi(\Gamma,{\mathrm{Sel}}(A/F^{\mathrm{cyc}}))\prod_{v\in S^{\prime}}\frac{\#A^{*}_{p^{\infty}}(F_{v})}{\#H^{1}(\Gamma_{w},A_{p^{\infty}}(F^{\mathrm{cyc}}_{w}))},$
where, for every $v\in S^{\prime}$, $w$ is a fixed prime of $F^{\mathrm{cyc}}$ above $v$.
The above result is a natural analogue of Zerbes [40], where the result was proved in the context of number fields. In the function field context, our result generalises previous computations of Sechi [31] and Valentino [36]. In the case $A$ is an elliptic curve, we may recover the formula of Sechi with Euler factors of the complex $L$-function appearing on the right hand side.
We conclude this introduction with an outline of the paper. In Section 2, we collect some basic notions concerning Iwasawa algebras that will be used throughout the paper. The proof of Theorem 1.3 will be given in Section 3, where the key observation is that the Galois group ${\mathrm{Gal}}(F_{S}/F^{\mathrm{cyc}})$, where $F_{S}$ is the maximal extension of $F$ contained in $\bar{F}$ which is unramified outside $S$, has $p$-cohomological dimension one (see Lemma 3.2). In Section 4, we establish the surjectivity of the localisation map defining the Selmer group ${\mathrm{Sel}}(A/F_{\infty})$ using Jannsen’s spectral sequence and Nekovář’s duality theorem. This will be a key input to our computation of Akashi series in Section 5. In Section 6, we prove a control theorem, showing in particular that the restriction map ${\mathrm{Sel}}(A/F)\to{\mathrm{Sel}}(A/F^{\mathrm{cyc}})^{\Gamma}$ has finite kernel and cokernel. This allows us to relate the order of vanishing at $T=0$ of the characteristic power series of ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ to the ${\mathbb{Z}}_{p}$-corank of the Selmer group ${\mathrm{Sel}}(A/F)$. In Section 7, we relate the order of vanishing of characteristic elements evaluated at Artin representations to the Selmer coranks and their twists in the intermediate subextensions of the $p$-adic Lie extension. This, in particular, gives us Theorem 1.4. We then give the definition of the generalised Euler characteristic in Section 8, following the work of Zerbes [40]. Finally, the computations of the generalised Euler characteristics will be carried out in Section 9 under Hypothesis $(\mathbf{H})$ to obtain Theorem 1.5.
2. Prerequisites on Iwasawa algebras
Throughout $p$ will denote a fixed prime. Let $G$ be a compact $p$-adic Lie group without $p$-torsion. The Iwasawa algebra of $G$ over ${\mathbb{Z}}_{p}$ is defined by
[…]
where $U$ runs over the open normal subgroups of $G$ and the inverse limit is taken with respect to the usual projection maps. Suppose for now that $G$ is pro-$p$. Then the ring $\Lambda(G)$ has a skew field of fractions $\mathcal{Q}(G)$ which allows us to define the rank of a $\Lambda(G)$-module $M$ via the following:
${\mathrm{rank}}_{\Lambda(G)}(M)=\dim_{\mathcal{Q}(G)}\mathcal{Q}(G)\otimes_{\Lambda(G)}M.$
The module $M$ is said to be torsion if ${\mathrm{rank}}_{\Lambda(G)}(M)=0$.
Now, suppose that $G$ is not necessarily pro-$p$. By Lazard’s theorem (see [12, Corollary 8.34]), one can always find an open subgroup $G^{\prime}$ of $G$ which is pro-$p$. Then a $\Lambda(G)$-module $M$ is said to be torsion over $\Lambda(G)$ if it is torsion over $\Lambda(G^{\prime})$ in the above sense. We shall also make use of the following equivalent formulation of a torsion $\Lambda(G)$-module: ${\mathrm{Hom}}_{\Lambda(G)}(M,\Lambda(G))=0$.
3. Selmer groups and the ${\mathfrak{M}}_{H}(G)$-Conjecture
Let $p$ be any prime number. Throughout, $F$ will always denote a global function field of characteristic coprime to $p$. We write $k$ for its constant field. Suppose that $A$ is an abelian variety defined over $F$. Let $F_{\infty}$ be a $p$-adic Lie extension of $F$ satisfying the following admissibility conditions:
  1. (i)
    $F_{\infty}$ contains the cyclotomic ${\mathbb{Z}}_{p}$-extension $F^{\mathrm{cyc}}$ of $F$.
  2. (ii)
    $G={\mathrm{Gal}}(F_{\infty}/F)$ is a compact $p$-adic group without $p$-torsion.
  3. (iii)
    $F_{\infty}/F$ is unramified outside of a finite set of primes.
In particular, $F^{\mathrm{cyc}}$ itself satisfies these conditions. Let $S$ be a nonempty finite set of primes of $F$ containing all the primes where the abelian variety $A$ has bad reduction, and all the primes of $F$ which ramify in $F_{\infty}$.
Let $\bar{F}$ denote a separable closure of $F$. Given any extension ${\mathfrak{K}}$ inside $\bar{F}$, we can define the $p$-primary Selmer group ${\mathrm{Sel}}(A/{\mathfrak{K}})$ of $E$ over ${\mathfrak{K}}$ using the Galois cohomology groups:
(3.1) ${\mathrm{Sel}}(A/{\mathfrak{K}})={\rm ker}\left(H^{1}({\mathfrak{K}},A_{p^{\infty}})\rightarrow\prod_{w}H^{1}({\mathfrak{K}}_{w},A)\right),$
where $A_{p^{\infty}}$ denotes the group of all $p$-power division points in $A(\bar{F})$. We note that when ${\mathfrak{K}}$ is an infinite extension of $F$, then ${\mathfrak{K}}_{w}$ is given as the union of completions at $w$ of all finite extensions over $F$ contained in ${\mathfrak{K}}$. If ${\mathfrak{K}}$ is Galois over $F$, the natural action of ${\mathrm{Gal}}({\mathfrak{K}}/F)$ on $H^{1}({\mathfrak{K}},A_{p^{\infty}})$ turns ${\mathrm{Sel}}(A/{\mathfrak{K}})$ into a ${\mathrm{Gal}}({\mathfrak{K}}/F)$-module.
Let $F_{S}$ be the maximal extension of $F$ contained in $\bar{F}$ which is unramified outside $S$. By the choice of $S$, we know $F_{\infty}$ is contained in $F_{S}$. For any extension $K$ of $F$ contained in $F_{S}$, we write $G_{S}(K)$ for the Galois group ${\mathrm{Gal}}(F_{S}/K)$. When $L$ is a finite extension of $F$, we define each prime $v$ of $F$
$J_{v}(A/L)=\bigoplus_{w\mid v}H^{1}(L_{w},A)(p),$
where $w$ runs over all primes of $L$ lying above $v$. Since $L_{w}$ has characteristic different from $p$, we have $H^{1}(L_{w},A)(p)\simeq H^{1}(L_{w},A_{p^{\infty}})$ by Kummer theory. We shall make this identification without any further mention.
Let $\bar{L}_{w}$ be a separable closure of $L_{w}$, by choosing an embedding $\bar{L}\subset\bar{L}_{w}$, we can view ${\mathrm{Gal}}(\bar{L}_{w}/L_{w})$ as a subgroup of ${\mathrm{Gal}}(\bar{L}/L)$. We then have the following localisation map
(3.2) $\lambda_{S}(A/L):H^{1}(G_{S}(L),A_{p^{\infty}})\to\bigoplus_{v\in S}J_{v}(A/L)$
induced by restriction.
For an infinite separable extension $K$ of $F$ contained in $F_{S}$, we define
$J_{v}(A/K)=\lim_{\longrightarrow}J_{v}(A/L),$
where $L$ runs over all finite extensions of $F$ contained in $K$ and the limit is taken with respect to the restriction maps. We similarly define the localisation map
$\lambda_{S}(A/K)=\lim_{\longrightarrow}\lambda_{S}(A/L).$
We will use the following result on Selmer groups in this paper, which can be proven in the same way as in the case of elliptic curves over number fields (see [5, Lemma 2.3]).
Proposition 3.1.
Let $K$ be an extension of $F$ contained in $F_{S}$. Then ${\mathrm{Sel}}(A/K)$ satisfies the exact sequence
$0\to{\mathrm{Sel}}(A/K)\to H^{1}(G_{S}(K),A_{p^{\infty}})\stackrel{{\scriptstyle\lambda_{S}(A/K)}}{{\longrightarrow}}\bigoplus_{v\in S}J_{v}(A/K).$
The remainder of the section will be devoted to proving the $\mathfrak{M}_{H}(G)$-conjecture for the Pontryagin dual $X(A/F_{\infty})$ of the $p$-primary Selmer group ${\mathrm{Sel}}(A/F_{\infty})$. In order to achieve this, we start with the following observation which will be crucial in our subsequent discussion.
Lemma 3.2.
We have
$\mathrm{cd}_{p}(G_{S}(F^{\mathrm{cyc}}))=1.$
Proof.
Since the set $S$ is nonempty and $\mathrm{char}(F)\neq p$, Theorem 10.1.2 (i)(c) from [28] tells us that $\mathrm{cd}_{p}(G_{S}(\bar{k}F))=1$. Note that $G_{S}(\bar{k}F)$ is a closed subgroup of $G_{S}(F^{\mathrm{cyc}})$, and its quotient ${\mathrm{Gal}}(\bar{k}F/F^{\mathrm{cyc}})$ is a profinite group, where every open subgroup has an index prime to $p$. Thus, by applying Proposition 3.3.5 (i) from [28], we find $\mathrm{cd}_{p}(G_{S}(F^{\mathrm{cyc}}))=\mathrm{cd}_{p}(G_{S}(\bar{k}F))$, and the lemma follows. ∎
Lemma 3.3.
The cohomology group $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})$ is cofinitely generated over $\Lambda(H)$. In particular, $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})^{\vee}$ is a torsion $\Lambda(G)$-module.
Proof.
Without loss of generality, we may assume that both $G$ and consequently $H$, are pro-$p$ without $p$-torsion. Therefore, by an application of Nakayama’s lemma, we are reduced to showing that $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})^{H}$ is cofinitely generated over ${\mathbb{Z}}_{p}$. Since $H^{2}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})=H^{2}(G_{S}(F_{\infty}),A_{p^{\infty}})=0$ by Lemma 3.2, the spectral sequence
$H^{i}\left(H,H^{j}(G_{S}(F_{\infty}),A_{p^{\infty}})\right)\Rightarrow H^{i+j}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})$
degenerates to yield an exact sequence
(3.3) $0\to H^{1}(H,A_{p^{\infty}}(F_{\infty}))\to H^{1}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})\to H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})^{H}\to H^{2}(H,A_{p^{\infty}}(F_{\infty}))\to 0$
[…]
$H^{1}(G_{S}(F_{\mathrm{cyc}}),A_{p})\twoheadrightarrow H^{1}(G_{S}(F_{\mathrm{cyc}}),A_{p^{\infty}})[p].$
Hence, it suffices to show $H^{1}(G_{S}(F_{\mathrm{cyc}}),A_{p})$ is finite, or equivalently, ${\mathrm{corank}}_{{\mathbb{F}}_{p}[[\Gamma]]}\left(H^{1}(G_{S}(F_{\mathrm{cyc}}),A_{p})\right)=0$. Since ${\mathrm{corank}}_{{\mathbb{F}}_{p}[[\Gamma]]}\left(H^{0}(G_{S}(F^{\mathrm{cyc}}),A_{p})\right)=0$ and $H^{2}(G_{S}(F^{{\mathrm{cyc}}}),A_{p})=0$, we have
(3.4) ${\mathrm{corank}}_{{\mathbb{F}}_{p}[[\Gamma]]}\left(H^{1}(G_{S}(F^{\mathrm{cyc}}),A_{p})\right)=\sum_{i=0}^{2}(-1)^{i+1}{\mathrm{corank}}_{{\mathbb{F}}_{p}[[\Gamma]]}\left(H^{i}(G_{S}(F^{\mathrm{cyc}}),A_{p})\right).$
It then follows from the spectral sequence
$H^{i}\left(\Gamma,H^{j}(G_{S}(F^{\mathrm{cyc}}),A_{p})\right)\Rightarrow H^{i+j}(G_{S}(F),A_{p})$
that the right hand side of (3.4) is
$\sum_{i=0}^{2}(-1)^{i+1}{\mathrm{corank}}_{{\mathbb{F}}_{p}}(H^{i}(G_{S}(F),A_{p})).$
The latter sum vanishes by a global Euler–Poincaré characteristic calculation (cf. [28, (8.7.4)]). This concludes the proof of the lemma. ∎
We can now prove the main result of this section.
Theorem 3.4.
Let $A$ be an abelian variety defined over a global function field $F$ of characteristic prime to $p$. Let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Then the Selmer group ${\mathrm{Sel}}(A/F_{\infty})$ of $A$ over $F_{\infty}$ is a cofinitely generated $\Lambda(H)$-module. In particular, the $\mathfrak{M}_{H}(G)$-conjecture holds for $X(A/F_{\infty})$.
Furthermore, ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ is a cofinitely generated ${\mathbb{Z}}_{p}$-module.
Proof.
Since ${\mathrm{Sel}}(A/F_{\infty})$ is contained in $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})$, the conclusion is now an immediate consequence of Lemma 3.3. ∎
4. Surjectivity of localisation map
[…]
$\mathrm{ker}\Big{(}H^{1}(G_{S}(L),T_{p}A)\to\prod_{w}H^{1}(L_{w},T_{p}A)\Big{)},$
where the product runs over all primes $w$ of $L$. We then define the compact Selmer group over $F_{\infty}$ by
[…]
Define the continuous Iwasawa cohomology of $T_{p}A$ by
[…]
where the inverse limit is taken over all finite Galois extensions $L$ of $F$ contained in $F_{\infty}$, via the corestriction maps. Note in particular that ${\mathfrak{S}}(A/F_{\infty})$ is a submodule of $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)$.
We now record the following result which will be a key ingredient for our eventual proof of the surjectivity.
Lemma 4.1.
$H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)$ is a torsion $\Lambda(G)$-module. Moreover, if $\dim(G)\geq 2$ or if $A_{p^{\infty}}(F_{\infty})$ is finite, then $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)=0$.
Proof.
We consider the following spectral sequence of Jannsen (see [16, Theorem 1, Theorem 11]; see also [23, Theorem 4.5.1] and [27, Theorem 8.5.6]))
$E_{2}^{i,j}={\mathrm{Ext}}^{i}_{\Lambda(G)}\left(H^{j}\left(G_{S}(F_{\infty}),A_{p^{\infty}})\right)^{\vee},\Lambda(G)\right)\Rightarrow H^{i+j}_{\mathrm{Iw}}(F_{\infty},T_{p}A).$
The low degree terms corresponding to this spectral sequence give us the following exact sequence
(4.1) $0\to{\mathrm{Ext}}^{1}_{\Lambda(G)}\left(A_{p^{\infty}}(F_{\infty})^{\vee},\Lambda(G)\right)\to H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)\to{\mathrm{Ext}}^{0}_{\Lambda(G)}\left(H^{1}\left(G_{S}(F_{\infty}),A_{p^{\infty}})\right)^{\vee},\Lambda(G)\right).$
By Lemma 3.3, we have ${\mathrm{rank}}_{\Lambda(G)}\left(H^{1}\left(G_{S}(F_{\infty}),A_{p^{\infty}})\right)^{\vee}\right)=0$, so the right-most term vanishes. Since ${\mathrm{Ext}}^{1}_{\Lambda(G)}\left(A_{p^{\infty}}(F_{\infty})^{\vee},\Lambda(G)\right)$ is a torsion $\Lambda(G)$-module (cf.  [37, Proposition 3.5 (iii) (a)]), it follows that $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)$ is a torsion $\Lambda(G)$-module. Now, $A_{p^{\infty}}(F_{\infty})^{\vee}$ is finitely generated over ${\mathbb{Z}}_{p}$, so the assumption $\dim(G)\geq 2$ tells us that $A_{p^{\infty}}(F_{\infty})^{\vee}$ is a pseudo-null module over $\Lambda(G)$, and the left-most term of (4.2) also vanishes. Thus, $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)=0$, as required. ∎
Theorem 4.2.
Let $A$ be an abelian variety defined over a global function field $F$ of characteristic prime to $p$. Let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Then the localisation map $\lambda_{S}(A/F_{\infty})$ is surjective, giving rise to the following short exact sequence:
$0\to{\mathrm{Sel}}(A/F_{\infty})\to H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\stackrel{{\scriptstyle\lambda_{S}(A/F_{\infty})}}{{\longrightarrow}}\bigoplus_{v\in S}J_{v}(A/F_{\infty})\to 0$
Proof.
As the conclusion of the theorem remains unchanged upon replacing the base field field $F$ by a larger field contained in $F_{\infty}$, we may assume that $G$ is pro-$p$. The Poitou-Tate sequence ([30, Proposition A.3.2]) gives an exact sequence
$0\to{\mathrm{Sel}}(A/F_{\infty})\to H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\stackrel{{\scriptstyle\lambda_{S}(A/F_{\infty})}}{{\longrightarrow}}\bigoplus_{v\in S}J_{v}(A/F_{\infty})\to{\mathfrak{S}}(A^{*}/F_{\infty})^{\vee}\to 0,$
where the final ``$0"$ term is a consequence of Lemma 3.2. It therefore remains to show that ${\mathfrak{S}}(A^{*}/F_{\infty})=0$.
We consider the following spectral sequence of Jannsen (see [16, Theorem 1, Theorem 11]; see also [23, Theorem 4.5.1] and [27, Theorem 8.5.6]))
$E_{2}^{i,j}={\mathrm{Ext}}^{i}_{\Lambda(G)}\left(H^{j}\left(G_{S}(F_{\infty}),A^{*}_{p^{\infty}})\right)^{\vee},\Lambda(G)\right)\Rightarrow H^{i+j}_{\mathrm{Iw}}(F_{\infty},T_{p}A^{*}).$
The low degree terms corresponding to this spectral sequence give us the following exact sequence
(4.2) $0\to{\mathrm{Ext}}^{1}_{\Lambda(G)}\left(A^{*}_{p^{\infty}}(F_{\infty})^{\vee},\Lambda(G)\right)\to H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A^{*})\to{\mathrm{Ext}}^{0}_{\Lambda(G)}\left(H^{1}\left(G_{S}(F_{\infty}),A^{*}_{p^{\infty}})\right)^{\vee},\Lambda(G)\right).$
By Lemma 3.3, we have ${\mathrm{rank}}_{\Lambda(G)}\left(H^{1}\left(G_{S}(F_{\infty}),A_{p^{\infty}})\right)^{\vee}\right)=0$, so the right-most term vanishes. Now, $A_{p^{\infty}}(F_{\infty})^{\vee}$ is finitely generated over ${\mathbb{Z}}_{p}$. Therefore, if $\dim(G)\geq 2$, then $A_{p^{\infty}}(F_{\infty})^{\vee}$ is a pseudo-null module over $\Lambda(G)$, and so the left-most term of (4.2) also vanishes. Thus, we have shown that ${\mathfrak{S}}(A^{*}/F_{\infty})=0$ when $\dim(G)\geq 2$.
It therefore remains to consider the case where $F_{\infty}=F^{\mathrm{cyc}}$. This case is slightly more involved, owing to the fact that we do not know the finiteness of $A_{p^{\infty}}(F^{\mathrm{cyc}})$ in general (but see Remark 4.3 in the case where $A$ is an elliptic curve without complex multiplication). By base changing $F$, if necessary, we may assume that every prime of $S$ is inert in $F^{\mathrm{cyc}}/F$. Thus, by abuse of notation, given $v\in S$, we also denote by $v$ the unique prime of $F^{\mathrm{cyc}}$ lying above $v$, and identify the Galois groups ${\mathrm{Gal}}(F^{\mathrm{cyc}}/F)$ with ${\mathrm{Gal}}(F^{\mathrm{cyc}}_{v}/F_{v})$, denoting them both by $\Gamma$. For each $v\in S$, we consider the following local version of Jannsen’s spectral sequence (see [16, Theorem 11], [23, Theorem 4.2.2] or [27, Theorem 5.2.6]) for the dual abelian variety $A^{*}$:
$E_{2}^{i,j}={\mathrm{Ext}}^{i}_{\Lambda(\Gamma)}\left(H^{j}\left(F^{\mathrm{cyc}}_{v},A_{p^{\infty}}^{*})\right)^{\vee},\Lambda(\Gamma)\right)\Rightarrow H^{i+j}_{\mathrm{Iw}}(F^{\mathrm{cyc}}_{v},T_{p}A^{*}).$
To simplify notation, we write $\mathrm{E}^{i}(-)={\mathrm{Ext}}^{i}_{\Lambda(\Gamma)}\left(-,\Lambda(\Gamma)\right)$. Analogously to the global case, the low degree terms of the above spectral sequence give an exact sequence
$0\to\mathrm{E}^{1}\left(A^{*}_{p^{\infty}}(F^{\mathrm{cyc}}_{v})^{\vee}\right)\to H^{1}_{\mathrm{Iw}}(F^{\mathrm{cyc}}_{v},T_{p}A^{*})\to\mathrm{E}^{0}\left(H^{1}(F^{\mathrm{cyc}}_{v},A^{*}_{p^{\infty}})^{\vee}\right).$
The global and local spectral sequences satisfy a functoriality property [23, Theorem 4.5.1], yielding the following commutative diagram:
[…]
where the vertical maps are induced by the natural localisation maps. Note that $\ker g={\mathfrak{S}}(A^{*}/F^{\mathrm{cyc}})$. Thus, it remains to show that $f$ is injective. Since $S$ is assumed to be non-empty, writing $f=\oplus_{v\in S}f_{v}$, it suffices to show that
$f_{v}:\mathrm{E}^{1}\left(A^{*}_{p^{\infty}}(F^{\mathrm{cyc}})^{\vee}\right)\to\mathrm{E}^{1}\left(A^{*}_{p^{\infty}}(F^{\mathrm{cyc}}_{v})^{\vee}\right)$
is injective for each $v\in S$. But since both $A^{*}_{p^{\infty}}(F^{\mathrm{cyc}})^{\vee}$ and $A^{*}_{p^{\infty}}(F^{\mathrm{cyc}}_{v})^{\vee}$ are finitely generated over ${\mathbb{Z}}_{p}$, it follows by [28, Corollary 5.5.7] that $f_{v}$ is given by
${\mathrm{Hom}}_{{\mathbb{Z}}_{p}}(A^{*}_{p^{\infty}}(F^{\mathrm{cyc}})^{\vee},{\mathbb{Z}}_{p})\to{\mathrm{Hom}}_{{\mathbb{Z}}_{p}}(A^{*}_{p^{\infty}}(F^{\mathrm{cyc}}_{v})^{\vee},{\mathbb{Z}}_{p}),$
which is injective since it is induced by the inclusion $A^{*}_{p^{\infty}}(F^{\mathrm{cyc}})\hookrightarrow A^{*}_{p^{\infty}}(F^{\mathrm{cyc}}_{v})$. This completes the proof of the theorem. ∎
Remark 4.3.
In the case where $A=E$ is an elliptic curve without complex multiplication, the proof of Theorem 4.2 can be simplified. By a theorem of Deuring, this is equivalent to saying that $E$ is not defined over a finite field; equivalently, $E$ is non-isotrivial. In particular, we do not have to separately consider the case $F_{\infty}=F^{\mathrm{cyc}}$, since Lemma 4.1 holds without the assumption $\dim(G)\geq 2$. One way to see this is the result of Ulmer, which says that $E(\bar{k}F)$ is a finitely generated abelian group (see [35, §6.3]). Since $F^{\rm cyc}$ is a subfield of $\bar{k}F$, it follows that $E(F^{\rm cyc})$ is a finitely generated abelian group, and thus $E_{p^{\infty}}(F^{\rm cyc})$ is finite. It follows that $E_{p^{\infty}}(F^{\rm cyc})$ is pseudo-null over $\Lambda(G)$, and we always have $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}E)$ injecting into the ${\mathrm{Ext}}^{0}$-term.
5. Akashi series
Denote by $\mathcal{Q}(\Gamma)$ the field of fractions of $\Lambda(\Gamma)$. To each $M$ in ${\mathfrak{M}}_{H}(G)$, we now attach a non-zero element $\mathrm{Ak}(M)$ of $\mathcal{Q}(\Gamma)$, which is uniquely determined up to multiplication by a unit in $\Lambda(\Gamma)$. It can be shown (see [8, Lemma 3.1]) that the homology groups
$H_{i}(H,M)\qquad\text{for }i\geq 0$
are finitely generated torsion $\Lambda(\Gamma)$-modules. We fix from now on a topological generator of $\Gamma$, and identify $\Lambda(\Gamma)$ with the formal power series ring ${\mathbb{Z}}_{p}[[T]]$ by mapping this topological generator to $1+T$. Let $f_{M,i}\in\Lambda(\Gamma)$ be a characteristic power series for $H_{i}(H,M)$, and define
(5.1) $\mathrm{Ak}(M)=\prod_{i\geq 0}f_{M,i}^{(-1)^{i}}.$
This product is finite because $H_{i}(H,M)=0$ for $i\geq\dim(H)$, and it is well defined up to multiplication by a unit in $\Lambda(\Gamma)$ because each $f_{M,i}$ is. We call $\mathrm{Ak}(M)$ the Akashi series of $M$. If $Y$ is a discrete $p$-primary $G$-module such that $M=Y^{\vee}\in{\mathfrak{M}}_{H}(G)$, then we call $\mathrm{Ak}(M)$ also the Akashi series of $Y$, and by abuse of notation, we also denote it by $\mathrm{Ak}(Y)$. Via the duality
$H_{i}(H,M)=H^{i}(H,Y)^{\vee},$
we will work with $H$-cohomology groups in our computation of Akashi series without further mention. It is indeed a generalisation of the notion of characteristic power series, because if $M$ is a finitely generated torsion $\Lambda(\Gamma)$-module (we can view $H=\{1\}$), then the Akashi series of $M$ is exactly the $\Lambda(\Gamma)$-characteristic power series of $M$. In this case, we simply denote $\mathrm{Ak}(M)$ by $f_{M}$, or $f_{Y}$, again by abuse of notation.
Lemma 5.1.
Given an exact sequence
(5.2) $0\to M_{1}\to M_{2}\to M_{3}\to 0$
of modules in the category ${\mathfrak{M}}_{H}(G)$, their Akashi series satisfy $\mathrm{Ak}(M_{2})=\mathrm{Ak}(M_{1})\cdot\mathrm{Ak}(M_{3})$.
Proof.
This follows from considering the long exact sequence of $H$-homology derived from (5.2) and applying the multiplicativity of $\Lambda(\Gamma)$-characteristic series in short exact sequences. ∎
Remark 5.2.
We note that the original definition of ${\mathfrak{M}}_{H}(G)$ in [9] differs from the definition in the subsequent paper [8] or in this paper. One consequence of this is that it is no longer true that if $M$ is a module in ${\mathfrak{M}}_{H}(G)$ then $\mathrm{Ak}(M)=\mathrm{Ak}(M/M(p))$, as shown in [9, Lemma 4.1]. This is because $M(p)$ need not be finitely generated over $\Lambda(H)$, and so its homology group may not be finitely generated over ${\mathbb{Z}}_{p}$. Therefore, even though its homology group is annihilated by some power of $p$, we cannot conclude that they are finite. In general, $\mathrm{Ak}(M(p))$ can be non-trivial. In case $G$ is pro-$p$ without $p$-torsion, it is given by $\mathrm{Ak}(M(p))=p^{\mu(M)}$ (see [1, 1.6]), where $\mu(M)$ denotes the $\mu$-invariant of $M$ defined in [37, 4].
By virtue of Theorem 4.2, we have a short exact sequence
$0\to{\mathrm{Sel}}(A/F_{\infty})\to H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\to\bigoplus_{v\in S}J_{v}(A/F_{\infty})\to 0$
Since ${\mathrm{Sel}}(A/F_{\infty})$ and $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})$ are cofinitely generated $\Lambda(H)$-modules by Lemma 3.3 and Theorem 3.4, so is $\bigoplus_{v\in S}J_{v}(A/F_{\infty})$. Therefore, we may apply Lemma 5.1 to conclude that
(5.3) $\mathrm{Ak}\Big{(}H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\Big{)}=\mathrm{Ak}\Big{(}{\mathrm{Sel}}(A/F_{\infty})\Big{)}\mathrm{Ak}\left(\bigoplus_{v\in S}J_{v}(A/F_{\infty})\right).$
Therefore, the determination of $\mathrm{Ak}\Big{(}{\mathrm{Sel}}(A/F_{\infty})\Big{)}$ is reduced to computing $\mathrm{Ak}\Big{(}H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\Big{)}$ and $\mathrm{Ak}\left(\bigoplus_{v\in S}J_{v}(A/F_{\infty})\right)$. These will be the content of the next two lemmas.
Lemma 5.3.
$\mathrm{Ak}\Big{(}H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\Big{)}=f_{H^{1}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})}\prod_{i\geq 1}\left(f_{H^{i}\left(H,A_{p^{\infty}}(F_{\infty})\right)}\right)^{(-1)^{i}}$.
Proof.
Since $H^{2}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})=H^{2}(G_{S}(F_{\infty}),A_{p^{\infty}})=0$ by Lemma 3.2, the spectral sequence
$H^{i}\left(H,H^{j}(G_{S}(F_{\infty}),A_{p^{\infty}})\right)\Rightarrow H^{i+j}(G_{S}(F_{\infty}),A_{p^{\infty}})$
degenerates to yield an exact sequence
(5.4) $0\to H^{1}(H,A_{p^{\infty}}(F_{\infty}))\to H^{1}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})\to H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})^{H}\to H^{2}(H,A_{p^{\infty}}(F_{\infty}))\to 0$
and isomorphisms
(5.5) $H^{i}\left(H,H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\right)\simeq H^{i+2}(H,A_{p^{\infty}}(F_{\infty}))$
for $i\geq 1$. The conclusion now follows from a direct calculation using (5.4) and (5.5). ∎
Since $\mathrm{char}(F_{v})\neq p$ for each $v\in S$, it follows from a theorem of Iwasawa [28, Theorem 7.5.3] that the decomposition subgroup $G_{w}$ of $w$ in $G$ has dimension at most $2$. Let $S^{\prime}$ denote the subset of $S$ consisting of primes whose inertia group in $G$ is infinite. We note that this is equivalent to requiring that the decomposition subgroup $G_{w}$ has dimension exactly $2$. Indeed, since the residue characteristic is prime to $p$ and $G$ is a $p$-adic Lie group, the wild inertia subgroup is finite. Moreover, the maximal unramified $p$-extension of $F_{v}$ is $F_{w}^{\mathrm{cyc}}$, and the maximal tamely ramified extension of $F_{v}$ is a topologically cyclic extension over the maximal unramified extension.
Lemma 5.4.
We have $\mathrm{Ak}\left(\bigoplus_{v\in S}J_{v}(A/F_{\infty})\right)=\prod_{v\in S\setminus S^{\prime}}f_{J_{v}(A/F^{\mathrm{cyc}})}$
Proof.
For each $v\in S$, we have
$J_{v}(A/F_{\infty})$ $=\bigoplus_{w\mid v}\mathrm{Ind}_{H_{w}}^{H}H^{1}(F_{\infty,w},A_{p^{\infty}}),$
where the sum runs over places $w$ of $F^{\mathrm{cyc}}$ lying above $v$, and we also write $w$ for a fixed prime in $F_{\infty}$ lying above $w$. We claim that $J_{v}(A/F_{\infty})=0$ for $v\in S^{\prime}$. Indeed, if we let $w$ denote a prime of $F_{\infty}$ above $v$, then $\dim G_{w}=2$ by our assumption that $v\in S^{\prime}$. The theorem of Iwasawa [28, Theorem 7.5.3] then tells us that $F_{\infty,w}$ has no nontrivial $p$-extension. Hence $H^{1}(F_{\infty,w},A_{p^{\infty}})=0$ and this establishes our claim.
Taking this claim into account, it remains to compute $\mathrm{Ak}\left(J_{v}(A/F_{\infty})\right)$ for $v\in S\setminus S^{\prime}$. For a prime $v$ of $F$, we write $w$ for a prime of $F^{\mathrm{cyc}}$ lying above $v$, and by abuse of notation, we also write $w$ for a fixed prime in $F_{\infty}$ lying above $w$. We then denote by $H_{w}$ the decomposition group in $H$. Shapiro’s lemma gives
(5.6) $H^{i}(H,J_{v}(A/F_{\infty}))\simeq\bigoplus_{w\mid v}H^{i}(H_{w},H^{1}(F_{\infty,w},A_{p^{\infty}})).$
If $v\in S\>\backslash\>S^{\prime}$, then $\mathrm{cd}_{p}(H_{w})=0$. Thus, the right hand side of (5.6) vanishes for all $i\geq 1$. For the $J_{v}(A/F_{\infty})^{H}$ term, one can check easily that the restriction map
$J_{v}(A/F^{\mathrm{cyc}})\rightarrow J_{v}(A/F_{\infty})^{H}$
is an isomorphism using the fact that $\mathrm{cd}_{p}(H_{w})=0$. In conclusion, we have
$\mathrm{Ak}\Big{(}J_{v}(A/F_{\infty})\Big{)}=\left\{\begin{array}[]{ll}f_{J_{v}(A/F^{\mathrm{cyc}})}&\hbox{if $v\in S\setminus S^{\prime}$;}\\ 0,&\hbox{if $v\in S^{\prime}$.}\end{array}\right.$
Theorem 5.5.
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$. Let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Then we have
$\mathrm{Ak}\left({\mathrm{Sel}}(A/F_{\infty})\right)=f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})}\prod_{i\geq 1}\left(f_{H^{i}\left(H,A_{p^{\infty}}(F_{\infty})\right)}\right)^{(-1)^{i}}\prod_{v\in S^{\prime}}f_{J_{v}(A/F^{\mathrm{cyc}})}.$
Proof.
Putting the calculations of Lemma 5.3 and 5.4 into (5.3), we obtain
$f_{H^{1}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})}\prod_{i\geq 1}\left(f_{H^{i}\left(H,A_{p^{\infty}}(F_{\infty})\right)}\right)^{(-1)^{i}}=\mathrm{Ak}\left({\mathrm{Sel}}(A/F_{\infty})\right)\prod_{v\in S\setminus S^{\prime}}f_{J_{v}(A/F^{\mathrm{cyc}})}.$
Thanks to Theorem 4.2, there is a short exact sequence
$0\to{\mathrm{Sel}}(A/F^{\mathrm{cyc}})\to H^{1}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})\to\bigoplus_{v\in S}J_{v}(A/F^{\mathrm{cyc}})\to 0,$
where all the terms are cofinitely generated over ${\mathbb{Z}}_{p}$ by Lemma 3.3 and Theorem 3.4. Consequently, we have
$f_{H^{1}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})}=f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})}\prod_{v\in S}f_{J_{v}(A/F^{\mathrm{cyc}})}.$
Combining all of the above, we obtain the required conclusion of the theorem. ∎
We record an interesting corollary of our Theorem 5.5 which is an analogue of a result of Howson [15] (also see [6]).
Corollary 5.6.
Retain the settings of Theorem 5.5. Suppose further that $G$ is pro-$p$ with dimension $\geq 2$. We then have the following formula
$\begin{split}\mathrm{rank}_{{\mathbb{Z}}_{p}[[H]]}X(A/F_{\infty})=\mathrm{rank}_{{\mathbb{Z}}_{p}}X(A/F^{\mathrm{cyc}})-\mathrm{rank}_{{\mathbb{Z}}_{p}}A_{p^{\infty}}(F^{\mathrm{cyc}})^{\vee}\\ +\sum_{w\in S^{\prime}(F^{\mathrm{cyc}})}\mathrm{rank}_{{\mathbb{Z}}_{p}}A_{p^{\infty}}(F^{{\mathrm{cyc}}}_{w})^{\vee}.\end{split}$
Proof.
In the discussion of the proof, whenever $N$ is a finitely torsion ${\mathbb{Z}}_{p}[[\Gamma]]$-module with trivial $\mu$-invariant, we shall represent the characteristic series $f_{N}$ by a distinguished polynomial, and so it makes sense to speak of $\deg f_{N}$. In particular, we have ${\mathrm{rank}}_{{\mathbb{Z}}_{p}}(N)=\deg f_{N}$. Now, if $M$ is a ${\mathbb{Z}}_{p}[[G]]$-module which is finitely generated over ${\mathbb{Z}}_{p}[[H]]$, then $H_{i}(H,M)$ is finitely generated over ${\mathbb{Z}}_{p}$ for every $i$. Hence, we have
${\mathrm{rank}}_{{\mathbb{Z}}_{p}[[H]]}(M)=\sum_{i\geq 0}(-1)^{i}{\mathrm{rank}}_{{\mathbb{Z}}_{p}}H_{i}(H,M)$
(cf. [15, Theorem 1.1]). Combining this with the above, we see that ${\mathrm{rank}}_{{\mathbb{Z}}_{p}[[H]]}(M)$ can be read off from $\mathrm{Ak}(M)$. More precisely, we have
${\mathrm{rank}}_{{\mathbb{Z}}_{p}[[H]]}(M)=\sum_{i\geq 0}(-1)^{i}\deg f_{H_{i}(H,M)}.$
Now, rewriting the conclusion of Theorem 5.5, we have
$\mathrm{Ak}\left({\mathrm{Sel}}(A/F_{\infty})\right)=\mathrm{Ak}\left(A_{p^{\infty}}(F_{\infty})\right)f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})}f_{A_{p^{\infty}}(F^{\mathrm{cyc}})}^{-1}\prod_{v\in S^{\prime}}f_{J_{v}(A/F^{\mathrm{cyc}})}.$
Since $\dim H\geq 1$ by hypothesis, we see that ${\mathrm{rank}}_{{\mathbb{Z}}_{p}[[H]]}(A_{p^{\infty}}(F_{\infty})^{\vee})=0$. It remains to show that
$\deg f_{J_{v}(A/F^{\mathrm{cyc}})}=\sum_{w|v}{\mathrm{rank}}_{{\mathbb{Z}}_{p}}A_{p^{\infty}}(F^{{\mathrm{cyc}}}_{w})^{\vee}$
for $v\in S^{\prime}$, where $w$ runs through the primes of $F^{\mathrm{cyc}}$ above $v$. For this, we note that $J_{v}(A/F^{\mathrm{cyc}})=\oplus_{w|v}H^{1}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}})$. Since $v\in S^{\prime}$, it follows from the proof of Lemma 5.4 that $H^{1}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}})=H^{1}(H_{w},A_{p^{\infty}}(F_{\infty,w}))$. Since $\dim H_{w}\geq 1$ and ${\mathrm{rank}}_{{\mathbb{Z}}_{p}[[H]]}(A_{p^{\infty}}(F_{\infty,w})^{\vee})=0$, we have
${\mathrm{rank}}_{{\mathbb{Z}}_{p}}H^{1}(H_{w},A_{p^{\infty}}(F_{\infty,w}))^{\vee}={\mathrm{rank}}_{{\mathbb{Z}}_{p}}H^{0}(H_{w},A_{p^{\infty}}(F_{\infty,w}))^{\vee}={\mathrm{rank}}_{{\mathbb{Z}}_{p}}A_{p^{\infty}}(F^{{\mathrm{cyc}}}_{w})^{\vee}.$
This completes the proof of the corollary. ∎
6. Control Theorem
In this section, we will prove the following control theorem which will allow us to relate the order of vanishing at $T=0$ of the characteristic power series of ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ to the ${\mathbb{Z}}_{p}$-corank of the Selmer group over ${\mathrm{Sel}}(A/F)$ over the base field $F$.
Theorem 6.1.
Let $A$ be an abelian variety defined over a global function field $F$ of characteristic different from $p$. Let $F^{\mathrm{cyc}}$ denote the cyclotomic ${\mathbb{Z}}_{p}$-extension of $F$. Write $\Gamma_{n}={\mathrm{Gal}}(F_{n}/F)$, where $F_{n}$ for the $n$-th layer with of the ${\mathbb{Z}}_{p}$-extension $F^{\mathrm{cyc}}/F$ satisfying $[F_{n}:F]=p^{n}$. Then the restriction map
${\mathrm{Sel}}(A/F_{n})\to{\mathrm{Sel}}(A/F^{\mathrm{cyc}})^{\Gamma_{n}}$
has finite kernel and cokernel which are bounded independently of $n$.
Proof.
We will closely follow the method of the proof of the control theorem for fine Selmer groups over a number field as presented in [20]. We consider the following commutative diagram with exact rows
[…]
Since $\Gamma_{n}$ has $p$-cohomological dimension $1$, the inflation-restriction sequence tells us that $\ker\beta_{n}=H^{1}(\Gamma_{n},A_{p^{\infty}}(F^{\mathrm{cyc}}))$ and $\mathrm{coker}~{}\!\beta_{n}=0$. Write $M=\left(A_{p^{\infty}}(F^{\mathrm{cyc}})\right)^{\vee}$. Then $M$ is a finitely generated ${\mathbb{Z}}_{p}$-module with a continuous action by $\Gamma$, so that it is a finitely generated torsion $\Lambda(\Gamma_{n})$-module for every $n$. It follows by [28, Proposition 5.3.20] that
${\mathrm{rank}}_{{\mathbb{Z}}_{p}}M_{\Gamma_{n}}={\mathrm{rank}}_{{\mathbb{Z}}_{p}}M^{\Gamma_{n}}<\infty.$
But we have ${\mathrm{rank}}_{{\mathbb{Z}}_{p}}M_{\Gamma_{n}}=0$ since $M_{\Gamma_{n}}=\left(A_{p^{\infty}}(F_{n})\right)^{\vee}$ is finite by the Mordell–Weil theorem. It follows that $M^{\Gamma_{n}}$ is also finite. Since $\Gamma_{n}$ has $p$-cohomological dimension $1$, we have
$H^{1}(\Gamma_{n},A_{p^{\infty}}(F^{\mathrm{cyc}}))=\left(A_{p^{\infty}}(F^{\mathrm{cyc}})\right)_{\Gamma_{n}},$
which is the Pontryagin dual of $M^{\Gamma_{n}}$. Since $M$ is a finitely generated ${\mathbb{Z}}_{p}$-module, the order of $M^{\Gamma_{n}}$ is bounded independently of $n$ by $M(p)$. It remains to show that $\ker\gamma_{n}$ is finite and bounded independently of $n$ in order to conclude that the same holds for $\ker\alpha_{n}$ and $\mathrm{coker}~{}\!\alpha_{n}$, by the snake lemma. For this, since $v\in S$ does not lie above $p$, it suffices to show that for each prime $w$ of $F^{\mathrm{cyc}}$ lying above $v\in S$, $H^{1}(\Gamma_{w},A_{p^{\infty}}(F^{\mathrm{cyc}}_{w}))$ is finite and bounded independently of $n$. This can be shown by replacing $\Gamma_{n}$ with $\Gamma_{w}$ and $M$ with $M_{w}=\left(A_{p^{\infty}}(F^{\mathrm{cyc}}_{w})\right)^{\vee}$, which is still a finitely generated ${\mathbb{Z}}_{p}$-module, in the above argument showing that $H^{1}(\Gamma_{n},A_{p^{\infty}}(F^{\mathrm{cyc}}))$ is finite and bounded independently of $n$. ∎
We have shown in Theorem 3.4 that $X(A/F^{\mathrm{cyc}})$ is a finitely generated torsion $\Lambda(\Gamma)$-module with $\mu$-invariant zero. Thus, the structure theorem for finitely generated $\Lambda(\Gamma)$-modules tells us that there is a pseudo-isomophism
(6.2) $X(A/F^{\mathrm{cyc}})\sim\bigoplus_{i=1}^{k}\Lambda(\Gamma)/(f_{i}^{a_{i}}),$
where we have identified $\Lambda(\Gamma)\simeq{\mathbb{Z}}_{p}[[T]]$ using our fixed topological generator of $\Gamma$, and the $f_{i}$ for $i=1,\ldots,k$ are polynomials of degree $\lambda_{i}$, say, which are distinguished and irreducible. Thus, $\Lambda(\Gamma)/(f_{i}^{a_{i}})$ is isomorphic to ${\mathbb{Z}}_{p}^{a_{i}\lambda_{i}}$ as abelian groups, and
${\mathrm{rank}}_{{\mathbb{Z}}_{p}}\left(X(A/F^{\mathrm{cyc}})\right)=\lambda=\sum_{i=1}^{k}a_{i}\lambda_{i}.$
Furthermore, we have $\prod_{i=1}^{k}f_{i}^{a_{i}}=f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})}$ up to a unit in $\Lambda(\Gamma)$.
We will make use of the following analogue of the semi-simplicity conjecture due to Greenberg, which was posed for elliptic curves over a number field, for abelian varieties over a global function field.
Conjecture 6.2 (Greenberg’s semi-simplicity conjecture for $A/F$).
Let $A$ be an abelian variety defined over a global function field $F$. Then the action of $\Gamma$ on $X(A/F^{\mathrm{cyc}})$ is completely reducible, that is, $a_{i}=1$ for all $i$ in (6.2).
We then have the following.
Proposition 6.3.
Let $A$ be an abelian variety defined over a global function field $F$. Then we have
${\mathrm{ord}}_{T=0}(f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})})\geq{\mathrm{corank}}_{{\mathbb{Z}}_{p}}({\mathrm{Sel}}(A/F)).$
In addition, equality holds when Conjecture 6.2 holds.
Proof.
We have shown in Theorem 6.1 that ${\mathrm{Sel}}(A/F)\to{\mathrm{Sel}}(A/F^{\mathrm{cyc}})^{\Gamma}$ has finite kernel and cokernel. Taking the Pontryagin dual, we see that the map $X(A/F^{\mathrm{cyc}})_{\Gamma}\to X(A/F)$ has finite kernel and cokernel. It follows that
${\mathrm{rank}}_{{\mathbb{Z}}_{p}}(X(A/F^{\mathrm{cyc}})_{\Gamma})={\mathrm{corank}}_{{\mathbb{Z}}_{p}}({\mathrm{Sel}}(A/F)).$
Now, we have ${\mathrm{rank}}_{{\mathbb{Z}}_{p}}(X(A/F^{\mathrm{cyc}})_{\Gamma})={\mathrm{rank}}_{{\mathbb{Z}}_{p}}\left(X(A/F^{\mathrm{cyc}})/TX(A/F^{\mathrm{cyc}})\right)$, and since ${\mathrm{rank}}_{{\mathbb{Z}}_{p}}\left(\Lambda(\Gamma)/(T)\right)=1$ and $\Lambda(\Gamma)/(f,T)\simeq{\mathbb{Z}}_{p}/f(0){\mathbb{Z}}_{p}$ for a distinguished polynomial $f$, the ${\mathbb{Z}}_{p}$-rank of the latter is $1$ if and only if $T\mid f$. It thus follows from (6.2) that ${\mathrm{rank}}_{{\mathbb{Z}}_{p}}\left(X(A/F^{\mathrm{cyc}})/TX(A/F^{\mathrm{cyc}})\right)$ is equal to the number of $f_{i}$ equal to $T$. Assuming Conjecture 6.2, this is precisely to the power of $T$ dividing $f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})}$, and the result follows. ∎
7. The order of vanishing of the characteristic element
In this section, we prove our main result on the order of vanishing of the characteristic element attached to the Selmer group over a $p$-adic Lie extension. We shall begin recalling the definition of these characteristic elements. As before, $G$ will denote a compact $p$-adic Lie group without $p$-torsion and $H$ is a closed normal subgroup of $G$ such that $\Gamma:=G/H\simeq{\mathbb{Z}}_{p}$. First, we introduce the characteristic element $\xi_{M}$ attached to $M\in\mathfrak{M}_{H}(G)$. In order to do this, we recall from [8] an equivalent definition of the category $\mathfrak{M}_{H}(G)$. A multiplicatively closed subset $\Sigma$ of $\Lambda(G)$ is said to be a left and right Ore set if, for each $f\in\Sigma$ and $x\in\Lambda(G)$, there exist $g_{1},g_{2}\in\Sigma$ and $y_{1},y_{2}\in\Lambda(G)$ such that
$fy_{1}=xg_{1}\quad\text{and}\quad y_{2}f=g_{2}x.$
We define
[…]
It is shown in [8, Theorem 2.4] that the set $\Sigma$ is a multiplicatively closed left and right Ore set in $\Lambda(G)$. Since $p\notin\Sigma$, we also need to consider
$\Sigma^{*}=\cup_{n\geq 0}p^{n}\Sigma,$
which is again a multiplicatively closed left and right Ore set in $\Lambda(G)$. We may therefore localise $\Lambda(G)$ at $\Sigma$ and at $\Sigma^{*}$ to obtain $\Lambda(G)_{\Sigma}$ and $\Lambda(G)_{\Sigma^{*}}=\Lambda(G)_{\Sigma}\left[\frac{1}{p}\right]$, respectively. It follows from [8, Proposition 2.3] that, given a finitely generated $\Lambda(G)$-module $M$, we have $M\in\mathfrak{M}_{H}(G)$ if and only if $M$ is $\Sigma^{*}$-torsion. Since the Iwasawa algebra $\Lambda(G)$ is noetherian and $G$ has no element of order $p$, $\Lambda(G)$ has finite global dimension equal to $\dim(G)+1$. Then $K_{0}(\Lambda(G))$ can be identified with the Grothendieck group of the category of all finitely generated $\Lambda(G)$-modules. It follows from the localisation sequence of $K$-theory for the Ore set $\Sigma^{*}$ that we have an exact sequence
$\cdots\to K_{1}(\Lambda(G))\to K_{1}(\Lambda(G)_{\Sigma^{*}})\stackrel{{\scriptstyle\partial_{G}}}{{\longrightarrow}}K_{0}(\mathfrak{M}_{H}(G))\to K_{0}(\Lambda(G))\to K_{0}(\Lambda(G)_{\Sigma^{*}})\to 0.$
By [8, Proposition 3.4], the map $\partial_{G}$ is surjective. In view of this result, we can define the notion of a characteristic element for each $M$ in $\mathfrak{M}_{H}(G)$.
Definition 7.1.
Given $M$ in $\mathfrak{M}_{H}(G)$, a characteristic element for $M$ is any element $\xi_{M}\in K_{1}(\Lambda(G)_{\Sigma^{*}})$ satisfying
$\partial_{G}(\xi_{M})=[M],$
where $[M]$ denotes the class of $M$ in $K_{0}(\Lambda(G))$.
We fix an algebraic closure ¯$\overline{{\mathbb{Q}}}_{p}$ of ${\mathbb{Q}}_{p}$. Let $\rho:G\to{\mathrm{GL}}_{n}(\mathcal{O})$ be an Artin representation, where $\mathcal{O}$ is the ring of integers of a finite extension of ${\mathbb{Q}}_{p}$ in ¯$\overline{{\mathbb{Q}}}_{p}$. We write $\Lambda_{\mathcal{O}}(\Gamma)$ for the Iwasawa algebra of $\Gamma$ with coefficients in $\mathcal{O}$, which we identify with the formal power series ring $\mathcal{O}[[T]]$ in an indeterminate $T$ by sending our fixed topological generator of $\Gamma$ to $1+T$. Let $\mathcal{Q}_{\mathcal{O}}(\Gamma)$ denote the field of fractions of $\Lambda_{\mathcal{O}}(\Gamma)$. If $M$ is a finitely generated $\Lambda(G)$-module, we put $M_{\mathcal{O}}=M\otimes_{{\mathbb{Z}}_{p}}\mathcal{O}$, and define
(7.1) $\mathrm{tw}_{\rho}(M)=M_{\mathcal{O}}\otimes_{\mathcal{O}}\mathcal{O}^{n}$
endowed with the diagonal action of $G$, where $G$ acts on $\mathcal{O}^{n}$ via $\rho$.
The representation $\rho$ defines a continuous group homomorphism
$G\to(M_{n}(\mathcal{O})\otimes_{{\mathbb{Z}}_{p}}\Lambda(\Gamma))^{\times}$
which maps $g$ to $\rho(g)\otimes\bar{g}$, where $\bar{g}$ denotes the image of $g$ in $\Gamma$. By [8, Lemma 3.3], this extends to a ring homomorphism
$\Lambda(G)_{\Sigma^{*}}\to M_{n}(\mathcal{Q}_{\mathcal{O}}(\Gamma)).$
From the functoriality of $K_{1}$-groups, this in turns induces a group homomorphism
$K_{1}(\Lambda(G)_{\Sigma^{*}})\to K_{1}(M_{n}(\mathcal{Q}_{\mathcal{O}}(\Gamma))).$
We then set $\Phi_{\rho}$ to be the composite map
(7.2) $\Phi_{\rho}:K_{1}(\Lambda(G)_{\Sigma^{*}})\to K_{1}(M_{n}(\mathcal{Q}_{\mathcal{O}}(\Gamma)))\stackrel{{\scriptstyle\text{Morita}}}{{\simeq}}K_{1}(\mathcal{Q}_{\mathcal{O}}(\Gamma))\simeq\mathcal{Q}_{\mathcal{O}}(\Gamma)^{\times}.$
Here, the first isomorphism is given by Morita’s equivalence. We also identify $\mathcal{Q}_{\mathcal{O}}(\Gamma)^{\times}$ with $\mathcal{Q}_{\mathcal{O}}(T)^{\times}$, where $\mathcal{Q}_{\mathcal{O}}(T)$ denotes the field of fractions of the power series ring $\mathcal{O}[[T]]$.
Definition 7.2 (Burns [3]).
Let $\xi\in K_{1}(\Lambda(G)_{\Sigma^{*}})$ and let $\rho$ be an Artin representation of $G$. Then there exists an integer $r_{\rho}(\xi)$ and $g(T)\in\mathcal{Q}_{\mathcal{O}}(T)^{\times}$ such that $g(0)\neq 0$ and
$\Phi_{\rho}(\xi)=T^{r_{\rho}(\xi)}g(T).$
We call $r_{\rho}(\xi)$ the order of vanishing of $\xi$ and denote by ${\mathrm{ord}}_{T=0}(\Phi_{\rho}(\xi))$.
We now relate the evaluation of the characteristic element of a module $M$ in the category $\mathfrak{M}_{H}(G)$ at Artin representations of $G$ to the Akashi series of the twists of $M$ by such representations defined by (7.1).
Lemma 7.3.
Let $M$ be a module in the category $\mathfrak{M}_{H}(G)$, and let $\xi_{M}\in K_{1}(\Lambda(G)_{\Sigma^{*}})$ denote its characteristic element. Then we have
$\Phi_{\rho}(\xi_{M})=\mathrm{Ak}(\mathrm{tw}_{\hat{\rho}}(M))\bmod\Lambda_{\mathcal{O}}(\Gamma)^{\times},$
where $\hat{\rho}$ denotes the contragradient representation of $\rho$, i.e. $\hat{\rho}(g)=\rho(g^{-1})^{t}$ for $g\in G$, where $t$ denotes the transpose matrix.
Proof.
By [8, Lemma 3.2], the module $\mathrm{tw}_{\hat{\rho}}(M)$ also lies in the category $\mathfrak{M}_{H}(G)$. Thus, $\mathrm{Ak}(\mathrm{tw}_{\hat{\rho}}(M))$ is well-defined, and the equality in the lemma is [8, (57)]. ∎
Given an Artin representation $\rho:G\to GL_{n}(\mathcal{O})$ of $G$ with coefficients in $\mathcal{O}$, let $F_{\rho}$ be the Galois extension of $F$ contained in $F_{\infty}$ such that ${\mathrm{Gal}}(F_{\infty}/F_{\rho})=\ker\rho$. We shall then write $G_{\rho}={\mathrm{Gal}}(F_{\infty}/F_{\rho})$. In the case the regular representation $\rho=\mathrm{reg}_{L}$ for a finite Galois extension $L$ of $F$ contained in $F_{\infty}$, we simply write $G_{L}={\mathrm{Gal}}(F_{\infty}/L)$ for $G_{\mathrm{reg}_{L}}$. We now introduce two hypotheses which will be useful for the computations in this section and in Section 9.
Hypothesis ($\mathbf{G}$). $H^{i}(G,A_{p^{\infty}}(F_{\infty}))$ is finite for all $i\geq 1$.
Hypothesis ($\mathbf{H}$). $H^{i}(H,A_{p^{\infty}}(F_{\infty}))$ is finite for all $i\geq 0$.
Remark 7.4.
Hypothesis $(\mathbf{G})$ is clearly satisfied if $A_{p^{\infty}}(F_{\infty})$ is finite. It also holds in the case when $F_{\infty}=F(A_{p^{\infty}})$ (see [2, Section 3]). Note that Hypothesis $(\mathbf{H})$ implies Hypothesis $(\mathbf{G})$.
Similarly, we write Hypothesis ($\mathbf{G_{\rho}}$) in case we replace $G$ with $G_{\rho}$, and Hypothesis $(\mathbf{G_{L}})$ in the case $\rho=\mathrm{reg}_{L}$ for a finite Galois extension $L$ of $F$ contained in $F_{\infty}$ with $G_{L}={\mathrm{Gal}}(F_{\infty}/L)$.
The next proposition provides a criterion for verifying Hypothesis $(\mathbf{H})$.
Proposition 7.5.
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$. Suppose that $A_{p^{\infty}}(F^{\mathrm{cyc}})$ is finite and $H\simeq{\mathbb{Z}}_{p}^{r}$. Then Hypothesis $(\mathbf{H})$ holds, i.e., $H^{i}(H,A_{p^{\infty}}(F_{\infty}))$ is finite for all $i\geq 0$.
Proof.
The conclusion of the proposition is a consequence of the following algebraic observation: If $H\cong{\mathbb{Z}}_{p}^{r}$, then for a given $\Lambda(H)$-module $M$, one has $H_{i}(H,M)$ being finite for every $i\geq 1$ whenever $H_{0}(H,M)$ is finite. This latter assertion essentially follows from [33, P. 56-57], although the discussion given there is rather terse. We refer the reader to [22, Proof of Theorem 2.3] for a detailed proof of this fact. ∎
We recall that, by Remark 4.3, the finiteness condition on $A_{p^{\infty}}(F^{\mathrm{cyc}})$ automatically holds when $A$ is an elliptic curve without complex multiplication.
We now prove the following proposition which illustrates the significance of Hypothesis $(\mathbf{G})$.
Proposition 7.6.
Let $A$ be an abelian variety defined over a global field of characteristic different from $p$. Suppose that Hypothesis $(\mathbf{G})$ holds. Then we have ${\mathrm{ord}}_{T=0}\left(f_{H^{i}(H,A_{p^{\infty}}(F_{\infty}))}\right)=0$ for all $i\geq 1$.
Proof.
To see that the characteristic series of $H^{i}(H,A_{p^{\infty}}(F_{\infty}))$ does not contribute a factor of $T$, it suffices to show that $H^{i}(H,A_{p^{\infty}}(F_{\infty}))^{\Gamma}$ is finite. Since $\mathrm{cd}_{p}\Gamma=1$, the spectral sequence
$H^{r}(\Gamma,H^{s}(H,A_{p^{\infty}}(F_{\infty})))\Rightarrow H^{r+s}(G,A_{p^{\infty}}(F_{\infty}))$
degenerates yielding a short exact sequence
$0\to H^{1}(\Gamma,H^{i-1}(H,A_{p^{\infty}}(F_{\infty})))\to H^{i}(G,A_{p^{\infty}}(F_{\infty}))\to H^{i}(H,A_{p^{\infty}}(F_{\infty}))^{\Gamma}\to 0$
for every $i\geq 1$. Since $H^{i}(G,A_{p^{\infty}}(F_{\infty}))$ is finite for every $i\geq 1$ by Hypothesis $(\mathbf{G})$, it follows that $H^{i}(H,A_{p^{\infty}}(F_{\infty}))^{\Gamma}$ is also finite for every $i\geq 1$, as required. ∎
Now, if $L$ is a finite Galois extension of $F$ contained in $F_{\infty}$, we write $\mathrm{reg}_{L}$ for the regular representation of ${\mathrm{Gal}}(L/F)$ and let $G_{L}$ denote the Galois group ${\mathrm{Gal}}(F_{\infty}/L)$.
Theorem 7.7.
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Denote by $\xi_{A}=\xi_{X(A/F_{\infty})}$ the characteristic element of $X(A/F_{\infty})$. Suppose that Hypothesis $(\mathbf{G_{L}})$ holds. Then we have
${\mathrm{ord}}_{T=0}(\Phi_{\mathrm{reg}_{L}}(\xi_{A}))\geq{\mathrm{corank}}_{{\mathbb{Z}}_{p}}({\mathrm{Sel}}(A/L)).$
In addition, equality holds when Greenberg’s semi-simplicity conjecture 6.2 holds for $A/L$.
Proof.
Set $H_{L}=G_{L}\cap H$ and $\Gamma_{L}={\mathrm{Gal}}(L^{\mathrm{cyc}}/L)$. Fix an appropriate power of the generator of $\Gamma$ so that it is a generator of $\Gamma_{L}$. With these choices of generators, we have
${\mathbb{Z}}_{p}[[T_{L}]]\simeq\Lambda(\Gamma_{L})\subseteq\Lambda(\Gamma)\simeq{\mathbb{Z}}_{p}[[T]].$
By the discussion in [21, Subsection 2.2 and Lemma 2.5], there is a restriction map
$\mathrm{res}:K_{1}(\Lambda(G)_{\Sigma^{*}})\longrightarrow K_{1}(\Lambda(G_{L})_{\Sigma_{G_{L}}^{*}})$
such that $\mathrm{res}(\xi_{A})$ is a characteristic element of $X(A/F_{\infty})$ viewed in $\mathfrak{M}_{H_{L}}(G_{L})$. On the other hand, [21, Proposition 2.10] tells us that
${\mathrm{ord}}_{T=0}(\Phi_{\mathrm{reg}_{L}}(\xi_{A}))={\mathrm{ord}}_{T_{L}=0}(\Phi_{\mathrm{reg}_{L},G_{L}}(\xi_{A})).$
Hence, in view of the above observations, it suffices to consider the case $F=L$. It follows from Lemma 7.3 that
$\Phi_{\mathrm{reg}_{F}}(\xi_{A})=\mathrm{Ak}({\mathrm{Sel}}(A/F_{\infty}))u(T)$
for some $u(T)\in\Lambda(T)^{\times}$. Taking the hypothesis of the theorem into account, we may apply Proposition 7.6 to conclude that ${\mathrm{ord}}_{T=0}\left(f_{H^{j}(H,A_{p^{\infty}}(F_{\infty}))}\right)=0$ for all $j\geq 1$. Combining this latter observation into Theorem 5.5, we obtain
${\mathrm{ord}}_{T=0}\left(\mathrm{Ak}\left({\mathrm{Sel}}(A/F_{\infty})\right)\right)={\mathrm{ord}}_{T=0}\left(f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})}\prod_{v\in S^{\prime}}f_{J_{v}(A/F^{\mathrm{cyc}})}\right).$
By Proposition 6.3, the order of vanishing of $f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})}$ at $T=0$ is greater or equal to ${\mathrm{corank}}_{{\mathbb{Z}}_{p}}({\mathrm{Sel}}(A/F))$, where equality holds if Conjecture 6.2 is valid. Hence it remains to show that the terms $f_{J_{v}(A/F^{\mathrm{cyc}})}$ at $v\in S^{\prime}$ does not contribute to the order of vanishing. For this, it suffices to show that $J_{v}(A/F^{\mathrm{cyc}})^{\Gamma}$ is finite. Indeed, noting that $\Gamma_{w}\simeq\Gamma\simeq{\mathbb{Z}}_{p}$ has $p$-cohomological dimension one, the inflation–restriction sequence gives us a surjection
$H^{1}(F_{v},A_{p^{\infty}})\twoheadrightarrow J_{v}(A/F^{\mathrm{cyc}})^{\Gamma}.$
By Tate duality (cf. [26, Chapter I, Remark 3.6]), we see that $H^{1}(F_{v},A_{p^{\infty}})$ is dual to $A^{*}_{p^{\infty}}(F_{v})$, and the latter is a finite group by Mattuck’s theorem for local fields. This completes the proof of the theorem. ∎
We now present a corollary which relates the characteristic element to the $L$-function of $A$ in the intermediate subfields of $F_{\infty}/F$. For a finite extension $L$ of $F$ contained in $F_{\infty}$, we write $L_{S}(A/L,s)$ for the $L$-function of $A$ with Euler factors removed at the primes of $L$ lying above $S$. The following is our corollary.
Corollary 7.8.
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Suppose that for a finite Galois extension $L$ of $F$ contained in $F_{\infty}$ the following statements are valid.
  1. (1)
    The Tate–Shafarevich group $\hbox{\cyr X}(A/L)$ is finite.
  2. (2)
    Greenberg’s semi-simplicity conjecture 6.2 holds for $A/L$.
  3. (3)
    Hypothesis $(\mathbf{G_{L}})$ holds.
Then we have
${\mathrm{ord}}_{T=0}(\Phi_{\mathrm{reg}_{L}}(\xi_{A}))={\mathrm{ord}}_{s=1}L_{S}(A/L,s).$
Proof.
Under the finiteness of $\hbox{\cyr X}(A/L)$, we have ${\mathrm{corank}}_{{\mathbb{Z}}_{p}}({\mathrm{Sel}}(A/L))={\mathrm{rank}}_{\mathbb{Z}}A(L)$. On the other hand, under the finiteness of $\hbox{\cyr X}(A/L)$, the deep works of Artin–Tate [34], Milne [25] and Kato–Trihan [18] show that ${\mathrm{rank}}_{\mathbb{Z}}A(L)={\mathrm{ord}}_{s=1}L_{S}(A/L,s)$. The conclusion of the corollary then follows from combining these equalities with Theorem 7.7. ∎
It is now natural to ask whether similar results can be obtained for Artin representations which are not the regular representation. In order to do this, we introduce the following notation. Let $\rho:G\to GL_{n}(\mathcal{O})$ be an Artin representation of $G$ with coefficients in $\mathcal{O}$. Denote by $W_{\rho}$ a free $\mathcal{O}$-module of rank $n$ realizing $\rho$. Let $F_{\rho}$ be the Galois extension of $F$ contained in $F_{\infty}$ such that ${\mathrm{Gal}}(F_{\infty}/F_{\rho})=\ker\rho$. We shall then write $G_{\rho}={\mathrm{Gal}}(F_{\infty}/F_{\rho})$.
Proposition 7.9.
Let $A$ be an abelian variety defined over a global field of characteristic different from $p$. Suppose that Hypothesis $(\mathbf{G_{\rho}})$ holds. Then we have ${\mathrm{ord}}_{T=0}\left(f_{H^{i}(H,\mathrm{tw}_{\rho}(A_{p^{\infty}})(F_{\infty}))}\right)=0$ for all $i\geq 1$.
Proof.
Let $H_{\rho}=G_{\rho}\cap H$. Consider the following spectral sequence
$H^{r}(H/H_{\rho},H^{s}(H_{\rho},\mathrm{tw}_{\rho}(A_{p^{\infty}})(F_{\infty}))\Rightarrow H^{r+s}(H,\mathrm{tw}_{\rho}(A_{p^{\infty}})(F_{\infty})).$
Since $H/H_{\rho}$ is finite, it remains to show that the characteristic series of $H^{i}(H_{\rho},\mathrm{tw}_{\rho}(A_{p^{\infty}})(F_{\infty}))$ does not contribute a factor of $T$. By the definition of $H_{\rho}$, the group $H_{\rho}$ acts trivially on $W_{\rho}$, and so we have
(7.3) $H^{i}(H_{\rho},\mathrm{tw}_{\rho}(A_{p^{\infty}})(F_{\infty}))=H^{i}(H_{\rho},A_{p^{\infty}}(F_{\infty}))\otimes_{\mathcal{O}}W_{\rho}.$
Since Proposition 7.6 tells us that the characteristic series of $H^{i}(H_{\rho},A_{p^{\infty}}(F_{\infty}))$ does not contribute a factor of $T$, we see from (7.3) that the same conclusion holds for $H^{i}(H_{\rho},\mathrm{tw}_{\rho}(A_{p^{\infty}})(F_{\infty}))$. ∎
Given any infinite separable extension $K$ of $F$ contained in $F_{S}$, we can define the twisted Selmer group ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/K)$ to be
${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/K)=\ker\left(H^{1}(G_{S}(K),\mathrm{tw}_{\rho}(A_{p^{\infty}}))\longrightarrow\bigoplus_{v\in S}J_{v}(\mathrm{tw}_{\rho}(A)/K)\right).$
Here, for each $v\in S$,
[…]
where $L$ runs over all finite extensions of $F$ contained in $K$, $w$ runs over all primes of $L$ lying above $v$ and the limit is taken with respect to the restriction maps. We shall write $X(\mathrm{tw}_{\rho}(A)/K)$ for the Pontryagin dual of ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/K)$.
Lemma 7.10.
We have the following short exact sequence
$0\longrightarrow{\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F_{\infty})\longrightarrow H^{1}(G_{S}(F_{\infty}),\mathrm{tw}_{\rho}(A_{p^{\infty}}))\longrightarrow\bigoplus_{v\in S}J_{v}(\mathrm{tw}_{\rho}(A)/F_{\infty})\longrightarrow 0$
of cofinitely generated $\Lambda_{\mathcal{O}}(H)$-modules.
Proof.
By the definition of $\rho$, we see that $G_{S}(F_{\infty})$ acts trivially on $W_{\rho}$, and so we have ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F_{\infty})={\mathrm{Sel}}(A/F_{\infty})\otimes_{\mathcal{O}}W_{\rho}$. Since ${\mathrm{Sel}}(A/F_{\infty})$ is cofinitely generated over $\Lambda(H)$ by Lemma 3.3, it follows that ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F_{\infty})$ is cofinitely generated over $\Lambda_{\mathcal{O}}(H)$. Via similar arguments, we see that $H^{1}(G_{S}(F_{\infty}),\mathrm{tw}_{\rho}(A_{p^{\infty}}))$ and $\bigoplus_{v\in S}J_{v}(\mathrm{tw}_{\rho}(A)/F_{\infty})$ are cofinitely generated over $\Lambda_{\mathcal{O}}(H)$. Since $W_{\rho}$ is free over $\mathcal{O}$, the short exact sequence of the lemma can be obtained from Theorem 4.2 by tensoring the short exact sequence in the said theorem with $W_{\rho}$. ∎
By a similar argument to that in [10, Proposition 2.5], the restriction map
${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F^{\mathrm{cyc}})\rightarrow{\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F_{\infty})^{H}$
has kernel which is cofinitely generated over $\mathcal{O}$. It then follows from this and the preceding lemma that ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F^{\mathrm{cyc}})$ is cofinitely generated over $\mathcal{O}$. Therefore, we may apply the structure theory for $\Lambda_{\mathcal{O}}(\Gamma)$-modules to obtain the following pseudo-isomorphism
(7.4) $X(\mathrm{tw}_{\rho}(A)/F^{\mathrm{cyc}})\sim\bigoplus_{j=1}^{t}\Lambda_{\mathcal{O}}(\Gamma)/(f_{j}^{\beta_{j}})$
where the $f_{j}$’s are irreducible distinguished polynomials.
We can now state the semi-simplicity conjecture of Greenberg for Artin twists of abelian varieties over a global function field.
Conjecture 7.11 (Greenberg’s semi-simplicity conjecture for $\mathrm{tw}_{\rho}(A)/F$).
Let $A$ be an abelian variety defined over a global function field $F$. Then the action of $\Gamma$ on $X(\mathrm{tw}_{\rho}(A)/F^{\mathrm{cyc}})$ is completely reducible, that is, $\beta_{j}=1$ for all $j$ in (7.4).
We are now in position to present the following result on the order of vanishing of characteristic elements evaluated at Artin representations.
Theorem 7.12.
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Let $\rho$ be an irreducible Artin representation of $G={\mathrm{Gal}}(F_{\infty}/F)$ and let $\xi_{A}$ denote a characteristic element of $X(A/F_{\infty})$. Suppose that Hypothesis $(\mathbf{G_{\rho}})$ holds. Then we have
${\mathrm{ord}}_{T=0}(\Phi_{\rho}(\xi_{A}))\geq{\mathrm{corank}}_{\mathcal{O}}({\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F)).$
In addition, equality holds when Greenberg’s semi-simplicity conjecture 7.11 holds for $\mathrm{tw}_{\rho}(A)/F$.
Proof.
The arguments used to prove Theorem 7.7 carry over for the Artin twists, taking Proposition 7.9 and Lemma 7.10 into account, and noting that $X(\mathrm{tw}_{\rho}(A)/F_{\infty})=\mathrm{tw}_{\hat{\rho}}X(A/F_{\infty})$ (cf. [7, Lemma 3.4]). ∎
8. The Generalised Euler characteristics
In this section, we give the definition of the generalised Euler characteristic. The idea of defining a finite Euler characteristic $\chi(G,Y)$ for a discrete $p$-primary $G$-module $Y$ when the cohomology groups $H^{i}(G,Y)$ are infinite was first conceived by Coates–Schneider–Sujatha in the form of the truncated Euler characteristic [9]. Our presentation here will follow that of Zerbes [40].
Let $W$ be a discrete $p$-primary $\Gamma$-module. Then we have
$H^{0}(\Gamma,W)=W^{\Gamma}\quad H^{1}(\Gamma,W)=W_{\Gamma},$
where $W_{\Gamma}$ is the maximal quotient of $W$ on which $\Gamma$ acts trivially. Hence, we have a map
$\phi_{W}:H^{0}(\Gamma,W)\to H^{1}(\Gamma,W),\quad f\mapsto\textrm{ residue class of }f.$
For a discrete $p$-primary $G$-module $Y$, we define
$d^{0}:H^{0}(G,Y)=H^{0}(\Gamma,Y^{H})\longrightarrow H^{1}(\Gamma,Y^{H})\hookrightarrow H^{1}(G,Y),$
where the middle map is $\phi_{Y^{H}}$ and the last map is given by inflation. Similarly, for $i\geq 1$, define
$d^{i}:H^{i}(G,Y)\twoheadrightarrow H^{0}(\Gamma,H^{i}(H,Y))\longrightarrow H^{1}(\Gamma,H^{i}(H,Y))\hookrightarrow H^{i+1}(G,Y),$
where the first map is given by restriction and the middle map is equal to $\phi_{H^{i}(H,Y)}$. Note that the first map is surjective since ${\rm cd}_{p}(\Gamma)=1$. We define $d^{-1}$ to be the zero map.
All the above properties for the morphisms involving in the definition of the complex $(H^{i}(G,Y),d^{i})$ will become clear in the proof of the following proposition.
Proposition 8.1.
The sequence $(H^{\bullet}(G,Y),d^{\bullet})$ forms a cochain complex.
Proof.
We have the Hochschild–Serre spectral sequence is given by
$H^{r}(\Gamma,H^{s}(H,Y))\Rightarrow H^{r+s}(G,Y).$
Since ${\rm cd}_{p}(\Gamma)=1$, we have $H^{r}(\Gamma,H^{s}(H,Y))=0$ unless $r=0$ or $1$. Thus, there exists a short exact sequence
(8.1) $0\to H^{1}(\Gamma,H^{s-1}(H,Y))\to H^{s}(G,Y)\to H^{0}(\Gamma,H^{s}(H,Y))\to 0$
for all $s\geq 1$. Now, we see that $d^{s}:H^{s}(G,Y)\to H^{s+1}(G,Y)$ is equal to the composition
$H^{s}(G,Y)\twoheadrightarrow H^{0}(\Gamma,H^{s}(H,Y))\stackrel{{\scriptstyle\phi_{H^{s}(H,X)}}}{{\longrightarrow}}H^{1}(\Gamma,H^{s}(H,Y))\hookrightarrow H^{s+1}(G,Y),$
where the surjective map and injective map are exactly given by the short exact sequence (8.1). Therefore, $d^{s+1}\circ d^{s}$ must factor through the following composition
$H^{1}(\Gamma,H^{s}(H,Y))\to H^{s+1}(G,Y)\to H^{0}(\Gamma,H^{s+1}(H,Y)),$
which is the exact sequence (8.1), and so must be the zero map. Hence, we have $d^{s+1}\circ d^{s}=0$, as desired. ∎
Definition 8.2.
Denote by ${\mathfrak{H}}^{\bullet}$ the cohomology of the complex in Proposition 8.1. If ${\mathfrak{H}}^{i}$ is finite for all $i\geq 0$, we say that the $G$-module $Y$ has finite generalised $G$-Euler characteristic $\chi(G,Y)$, defined by
$\chi(G,Y)=\prod_{i\geq 0}\#({\mathfrak{H}}^{i})^{(-1)^{i}}.$
Lemma 8.3.
The generalised $G$-Euler characteristic satisfies the following properties:
  1. (1)
    If the groups $H^{i}(G,Y)$ are finite for all $i$, the generalised $G$-Euler characteristic for $Y$ agrees with the usual $G$-Euler characteristic for $Y$.
  2. (2)
    If $H^{0}(G,Y)$ and $H^{1}(G,Y)$ are infinite and $H^{i}(G,Y)$ is finite for all $i\geq 2$. The module $Y$ has finite generalised $G$-Euler characteristic if and only if the map $d^{0}$ has finite kernel and cokernel, in which case
    (8.2) $\chi(G,Y)=\frac{\#(\ker(d^{0}))}{\#(\mathrm{coker}~{}\!(d^{0}))}\times\prod_{i\geq 2}\#\left(H^{i}(G,Y)\right)^{(-1)^{i}}.$
Proof.
This is essentially established in the work of Zerbes [40] but not explicitly stated in the form here. For the convenience of the readers, we shall supply a proof of the lemma. For the first assertion, it suffices to show that for a cochain complex $({\mathfrak{X}}^{\bullet},\delta^{\bullet})$ with ${\mathfrak{X}}^{i}$ finite for all $i\geq 0$, we have
$\prod^{\infty}_{i=0}\#(H^{i}({\mathfrak{X}}^{\cdot}))^{(-1)^{i}}=\prod^{\infty}_{i=0}\#({\mathfrak{X}}^{i})^{(-1)^{i}}.$
To see this, consider the following two short exact sequences
$0\to Z^{i}\to{\mathfrak{X}}^{i}\to B^{i+1}\to 0,$
$0\to B^{i}\to Z^{i}\to H^{i}({\mathfrak{X}}^{\cdot})\to 0$
As all modules in the sequences are finite, we have
#$\#({\mathfrak{X}}^{i})=\#(Z^{i})\cdot\#(B^{i+1}).$
Therefore, noting that #$\#(B^{0})=1$, we have
$\prod_{i\geq 0}\#({\mathfrak{X}}^{i})^{(-1)^{i}}=\prod_{i\geq 0}(\#(Z^{i})\cdot\#(B^{i+1}))^{(-1)^{i}}=\prod_{i\geq 0}\#(H^{i}({\mathfrak{X}}^{\cdot}))^{(-1)^{i}},$
and (1) follows. To prove the second assertion, we first note that $\ker(d^{0})={\mathfrak{H}}^{0}$. Thus, if $d^{0}$ has finite kernel and cokernel, then ${\mathfrak{H}}^{0}$ is finite. Now, from the exact sequence
$0\to{\mathfrak{H}}^{1}\to\mathrm{coker}~{}\!d^{0}\to H^{2}(G,Y),$
where the third map is induced by $d^{1}$, we see that ${\mathfrak{H}}^{1}$ is also finite. As ${\mathfrak{H}}^{i}$ are finite for $i\geq 2$ by assumption, $Y$ has finite generalised $G$-Euler characteristic. Conversely, suppose that $Y$ has finite generalised $G$-Euler characteristic. Since $\ker(d^{0})={\mathfrak{H}}^{0}$, we see that $d^{0}$ has finite kernel. Since $H^{2}(G,Y)$ is finite, $\ker(d^{2})$ is finite. Now, considering the exact sequence
$0\to{\mathfrak{H}}^{1}\to\mathrm{coker}~{}\!d^{0}\to\ker d^{2}\to{\mathfrak{H}}^{2}\to 0,$
we see that $\mathrm{coker}~{}\!(d^{0})$ is finite. It remains to show the formula (8.2). The complex $(H^{\bullet}(G,Y),d^{\bullet})$ induces the following cochain complex
$0\to\mathrm{coker}~{}\!(d^{0})\stackrel{{\scriptstyle\bar{d}^{1}}}{{\longrightarrow}}H^{2}(G,Y)\stackrel{{\scriptstyle d^{2}}}{{\longrightarrow}}\cdots.$
Since all terms in the cochain complex are finite, noting that $\ker(\bar{d}^{1})\simeq{\mathfrak{H}}_{1}$, we see from (1) that
(8.3) $\prod^{\infty}_{i=1}\#({\mathfrak{H}}^{i})^{(-1)^{i+1}}=\#(\mathrm{coker}~{}\!(d^{0}))\times\prod^{\infty}_{i=2}\#\left(H^{i}(G,Y)\right)^{(-1)^{i+1}}.$
Then noting that #$\#({\mathfrak{H}}^{0})=\#(\ker(d^{0}))$ and dividing both sides by the equality (8.3), we arrive at the following equality
$\chi(G,Y)=\frac{\#(\ker(d^{0}))}{\#(\mathrm{coker}~{}\!(d^{0}))}\times\prod_{i\geq 2}\#\left(H^{i}(G,Y)\right)^{(-1)^{i}},$
which is (8.2), as required. ∎
We now record the following important result which relates the generalised Euler characteristic to Akashi series.
Proposition 8.4.
Let $Y$ be a discrete $p$-primary $G$-module such that $Y^{\vee}\in{\mathfrak{M}}_{H}(G)$. If $Y$ has finite generalized $G$-Euler characteristic, then the leading term of its Akashi series $\mathrm{Ak}(Y)$ is $\alpha_{Y}T^{k}$, where $k$ is equal to the alternating sum
$k=\sum_{i\geq 0}(-1)^{i}{\rm corank}_{{\mathbb{Z}}_{p}}H^{i}(H,Y)^{\Gamma}$
and
$\chi(G,Y)=|\alpha_{Y}|^{-1}_{p}.$
Proof.
See [40, Proposition 2.10]. ∎
We note that it is in general not true that the generalised $G$-Euler characteristic is multiplicative in short exact sequences (see [40, Remark after Lemma 2.13]). However, the following lemma suffices for most of our purposes.
Lemma 8.5.
Let $0\to B_{1}\to B_{2}\to B_{3}\to B_{4}\to 0$ be an exact sequence of discrete $G$-modules with the property that $B_{1}$ and $B_{4}$ have finite $G$-Euler characteristics. Then $B_{2}$ has finite generalised $G$-Euler characteristic if and only if $B_{3}$ does.
Proof.
Set $C=\ker(B_{3}\to B_{4})$. By separately considering the following two short exact sequences
$0\to B_{1}\to B_{2}\to C\to 0$
$0\to C\to B_{3}\to B_{4}\to 0,$
we may assume that either $B_{1}=0$ or $B_{4}=0$. We shall prove the case when $B_{4}=0$, the case of $B_{1}=0$ being similar. Thus, we may suppose that we have a short exact sequence of $G$-modules
$0\to B_{1}\to B_{2}\to B_{3}\to 0$
with $B_{1}$ having finite $G$-Euler characteristics. In particular, we have that $H^{i}(G,B_{1})$ is finite for every $i\geq 0$. Taking the long exact sequence of $G$-cohomology of the short exact sequence above, we see that
$\psi_{i}:H^{i}(G,B_{2})\to H^{i}(G,B_{3})$
have finite kernel and cokernel for all $i\geq 0$. By the naturalness of the Hochschild–Serre spectral sequence and our definition of $d^{i}$, we see that $\psi_{i}$ are cochain maps between the complexes $(H^{\bullet}(G,B_{2}),d^{\bullet}_{B_{2}})$ and $(H^{\bullet}(G,B_{3}),d^{\bullet}_{B_{3}})$. Since the kernels and cokernels of these cochain map are finite, a direct homological algebraic argument shows that the induced maps on their cohomology also have finite kernel and cokernel. Hence it follows that $B_{2}$ has finite generalised Euler characteristic if and only if $B_{3}$ does. ∎
Remark 8.6.
One can even show that
$\chi(G,B_{2})\chi(G,B_{4})=\chi(G,B_{1})\chi(G,B_{3})$
in the preceding lemma. As this latter calculation is not required for the paper, we do not provide a detailed proof, leaving it to the interested readers.
9. Computations of the Generalised Euler Characteristics
In order to compute the Generalised Euler characteristics, we impose in this section Hypothesis $(\mathbf{H})$ on the abelian variety $A$ defined over a function field $F$, which we now recall:
Hypothesis $(\mathbf{H})$. $H^{i}(H,A_{p^{\infty}}(F_{\infty}))$ is finite for all $i\geq 0$.
Theorem 9.1.
Let $A$ be an abelian variety defined over a global field $F$ with $\mathrm{char}F\neq p$ and let $F_{\infty}$ be an admissible $p$-adic Lie extension of $F$. Suppose that Hypothesis $(\mathbf{H})$ holds. Then ${\mathrm{Sel}}(A/F_{\infty})$ has finite generalised $G$-Euler characteristic if and only if ${\mathrm{Sel}}(A/F^{\rm cyc})$ has finite generalised $\Gamma$-Euler characteristics. In the event of such, we have the equality
$\chi(G,{\mathrm{Sel}}(A/F_{\infty}))=\chi(\Gamma,{\mathrm{Sel}}(A/F^{\mathrm{cyc}}))\prod_{v\in S^{\prime}}\frac{\#A^{*}_{p^{\infty}}(F_{v})}{\#H^{1}(\Gamma_{w},A_{p^{\infty}}(F^{\mathrm{cyc}}_{w}))},$
where, for every $v\in S^{\prime}$, $w$ is a fixed prime of $F^{\mathrm{cyc}}$ above $v$.
Proof.
Consider the following fundamental commutative diagram
[…]
From the proof of Lemma 5.4, we see that the $p$-cohomological dimension of the decomposition group of $H$ at every prime of $F_{\infty}$ is at most one. It follows from this that the map $\gamma$ is surjective which in turn implies the same for $\phi$. The proof of the same lemma also gives $H^{i}(H,J_{v}(A/F_{\infty}))=0$ for every $i\geq 1$. Applying the snake lemma to the diagram (9.1), we obtain the following exact sequence
(9.2) $0\to\ker\alpha\to H^{1}(H,A_{p^{\infty}}(F_{\infty}))\to\ker\gamma\to\mathrm{coker}~{}\!\alpha\to H^{2}(H,A_{p^{\infty}}(F_{\infty}))\to 0$
and
$H^{i}(H,{\mathrm{Sel}}(A/F_{\infty}))\simeq H^{i+2}(H,A_{p^{\infty}}(F_{\infty}))$
for every $i\geq 1$. In view of this latter identification and Hypothesis $(\mathbf{H})$, we see that $H^{i}(H,{\mathrm{Sel}}(A/F_{\infty}))$ is finite for ech $i\geq 1$. It then follows from Lemma 8.3 that ${\mathrm{Sel}}(A/F_{\infty})$ has finite generalised $G$-Euler characteristic if and only if the map
$d^{0}:H^{0}(G,{\mathrm{Sel}}(A/F_{\infty}))=H^{0}(\Gamma,{\mathrm{Sel}}(A/F_{\infty})^{H})\stackrel{{\scriptstyle g}}{{\to}}H^{1}(\Gamma,{\mathrm{Sel}}(A/F_{\infty})^{H})\stackrel{{\scriptstyle h}}{{\hookrightarrow}}H^{1}(G,{\mathrm{Sel}}(A/F_{\infty}))$
has finite kernel and cokernel. Since $\mathrm{coker}~{}\!h=H^{1}(H,{\mathrm{Sel}}(A/F_{\infty}))^{\Gamma}\simeq H^{3}(H,A_{p^{\infty}}(F_{\infty}))^{\Gamma}$ is also finite, we see that the map $d_{0}$ has finite kernel and cokernel if and only if $g$ does. The latter is precisely equivalent to saying that ${\mathrm{Sel}}(A/F_{\infty})^{H}$ has finite generalised $\Gamma$-Euler characteristics.
On the other hand, from diagram (9.1), we have the following exact sequence
(9.3) $0\to\ker\alpha\to{\mathrm{Sel}}(A/F^{\rm cyc})\to{\mathrm{Sel}}(A/F_{\infty})^{H}\to\mathrm{coker}~{}\!\alpha\to 0$
Since $\ker\alpha\subseteq H^{1}(H,A_{p^{\infty}}(F_{\infty}))$, we see that $\ker\alpha$ is finite. We claim that $\mathrm{coker}~{}\!\alpha$ has finite $\Gamma$-Euler characteristic. Suppose for now the claim holds. Then it follows from Lemma 8.5 and (9.3) that ${\mathrm{Sel}}(A/F_{\infty})^{H}$ has finite generalised $\Gamma$-Euler characteristics if and only if ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ has finite generalised $\Gamma$-Euler characteristics, establishing the first assertion of the theorem. It therefore remains to prove our claim. Taking Hypothesis $(\mathbf{H})$ into account, the exact sequence (9.2) reduces our task to showing that $\ker\gamma$ has finite $\Gamma$-Euler characteristic. For this, we note that as observed in the proof of Theorem 7.7, there is a surjection
$H^{1}(F_{v},A_{p^{\infty}})\twoheadrightarrow J_{v}(A/F^{\mathrm{cyc}})^{\Gamma}.$
with $H^{1}(F_{v},A_{p^{\infty}})$ being finite. In particular, $J_{v}(A/F^{\mathrm{cyc}})^{\Gamma}$ is finite, and this in turn implies that $J_{v}(A/F^{\mathrm{cyc}})$ has a finite $\Gamma$-Euler characteristic.
From now on, we suppose that both generalised Euler characteristics in question are well-defined. Then we may apply Theorem 5.5 and Proposition 8.4 to conclude that
(9.4) $\chi(G,{\mathrm{Sel}}(A/F_{\infty}))=\chi(\Gamma,{\mathrm{Sel}}(A/F^{\mathrm{cyc}}))\prod_{v\in S^{\prime}}\chi(\Gamma,J_{v}(A/F^{\mathrm{cyc}})).$
Therefore, the final assertion of the theorem will follow once we can show that
(9.5) $\chi(\Gamma,J_{v}(A/F^{\mathrm{cyc}}))=\frac{\#A^{*}_{p^{\infty}}(F_{v})}{\#H^{1}(\Gamma_{w},A_{p^{\infty}}(F^{{\mathrm{cyc}}}_{w}))}$
for each $v$. The restriction map yields a short exact sequence
$0\to H^{1}(\Gamma_{w},A_{p^{\infty}}(F^{{\mathrm{cyc}}}_{w}))\to H^{1}(F_{v},A_{p^{\infty}})\to J_{v}(A/F^{\mathrm{cyc}})^{\Gamma}\to 0,$
where Tate duality tells us that $H^{1}(F_{v},A_{p^{\infty}})$ is isomorphic to the Pontryagin dual of $A^{*}_{p^{\infty}}(F_{v})$. Thus, we have
(9.6) #$\#J_{v}(A/F^{\mathrm{cyc}})^{\Gamma}=\frac{\#A^{*}_{p^{\infty}}(F_{v})}{\#H^{1}(\Gamma_{w},A_{p^{\infty}}(F^{{\mathrm{cyc}}}_{w}))}.$
On the other hand, we have
(9.7) $H^{1}\left(\Gamma,J_{v}(A/F^{\mathrm{cyc}})\right)\simeq H^{1}(\Gamma_{w},H^{1}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}}))$
by Shapiro’s lemma. Now, consider the following spectral sequence
$H^{r}(\Gamma_{w},H^{s}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}}))\Rightarrow H^{r+s}(F_{v},A_{p^{\infty}}).$
Since $\mathrm{cd}_{p}~{}\Gamma=1$, $H^{2}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}})=0$ (cf. [28, Theorem 7.1.8(i)]) and $H^{2}(F_{v},A_{p^{\infty}})=0$ (see [11, Lemma 1.12]), the spectral sequence degenerates yielding
(9.8) $H^{1}(\Gamma_{w},H^{1}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}}))=0.$
Combining (9.6), (9.7) and (9.8), we obtain (9.5), as required. This completes the proof of our theorem. ∎
Remark 9.2.
When $A=E$ is an elliptic curve, it follows from well-known properties of elliptic curves that
$\chi(\Gamma,J_{v}(E/F^{\rm cyc}))=|L_{v}(E,1)|_{p}$
where $L_{v}(E,1)$ denotes the Euler factor at $v$ of the complex $L$-function of $E$ over $F$, and $|\cdot|_{p}$ is the $p$-adic valuation of ${\mathbb{Q}}$, normalised so that $|p|_{p}=p^{-1}$. This recovers the formula of Zerbes in the number field setting and a formula of Sechi using the standard Euler characteristics in the function field setting.
Corollary 9.3.
Retain the settings of Theorem 9.1. Suppose further that Greenberg’s conjecture semi-simplicity conjecture 6.2 holds for $A/F$. Set $s_{A}=\mathrm{corank}_{{\mathbb{Z}}_{p}}{\mathrm{Sel}}(A/F)$. We then have
$\frac{1}{T^{s_{A}}}\Phi_{\mathrm{reg}_{F}}(\xi_{A})\Big{|}_{T=0}=\chi(\Gamma,{\mathrm{Sel}}(A/F^{\mathrm{cyc}}))\prod_{v\in S^{\prime}}\frac{\#A^{*}_{p^{\infty}}(F_{v})}{\#H^{1}(\Gamma_{w},A_{p^{\infty}}(F^{\mathrm{cyc}}_{w}))}.$
where, for every $v\in S^{\prime}$, $w$ is a fixed prime of $F^{\mathrm{cyc}}$ above $v$.
Proof.
This follows from a combination of Theorems 7.7 and 9.1. ∎